### v1

# Reasoning Grasping via Multimodal Large Language Model

Abstract
Despite significant progress in robotic systems for operation within human-centric environments, existing models still heavily rely on explicit human commands to identify and manipulate specific objects. This limits their effectiveness in environments where understanding and acting on implicit human intentions are crucial. In this study, we introduce a novel task: reasoning grasping, where robots need to generate grasp poses based on indirect verbal instructions or intentions. To accomplish this, we propose an end-to-end reasoning grasping model that integrates a multi-modal Large Language Model (LLM) with a vision-based robotic grasping framework. In addition, we present the first reasoning grasping benchmark dataset generated from the GraspNet-1 billion, incorporating implicit instructions for object-level and part-level grasping, and this dataset will soon be available for public access. Our results show that directly integrating CLIP or LLaVA with the grasp detection model performs poorly on the challenging reasoning grasping tasks, while our proposed model demonstrates significantly enhanced performance both in the reasoning grasping benchmark and real-world experiments.
I Introduction
Robotic grasping has long been a subject of extensive study. The utilization of CNN-based neural networks has demonstrated efficiency in generating high-quality grasping poses from visual input [32, 22, 19, 24, 34, 28]. One significant limitation of these methods is the lack of scene understanding, resulting in robots being unable to identify the objects they are grasping. Recent developments have introduced methods that interpret language instructions such as “I want a stapler” [7] or “Pick the food box in front of the ball” [43] for targeted grasping, enabling robots to pick objects as directed. However, those methods will struggle when dealing with implicit instructions, where the desired object is not explicitly named. For example, if a user says “I need to drink water”, an intelligent assistant robot should autonomously identify and grasp a cup from a selection of household items on the table, although “cup” does not appear in the user’s instruction. Such scenarios, prevalent in real-world setting, underscore the importance of advanced understanding and autonomy in robotic grasping — an area that remains a significant challenge.
Recently, the advance of large language models like ChatGPT [33] has introduced new possibilities for language reasoning, significantly enhancing our ability to process and interpret complex language structures. Beyond text, numerous studies have leveraged multi-modal Large Language Models (LLMs) for enhanced visual understanding and interaction [51, 12, 25], enabling these models to interpret and generate responses based on a combination of textual and visual inputs. However, these approaches have been confined largely to visual comprehension and text generation, lacking the capability to generate precise robotic actions. While there have been efforts to integrate LLMs with robotics [2, 17, 21, 37, 44, 18], these have predominantly focused on high-level planning or have modularized the LLM for reasoning while relying on traditional methods for downstream actions. Brohan et al. [4] utilize a Vision Language Model (VLM) to transfer web knowledge to robot control, but it requires extensive efforts to train on a large amount of data.
In this work, we introduce the novel task of Reasoning Grasping, where the robot is tasked with generating appropriate grasping poses based on implicit instructions. This task presents two significant challenges, primarily due to the nuanced requirements of interpreting complex, context-rich instructions and accurately generating the corresponding grasp poses.
To address these challenges, we propose a model that is uniquely designed to directly output grasp poses according to visual scenes and complex instructions in cluttered environments. Specifically, this model initially leverages a multi-modal LLM to interpret both images and instructions, producing a textual output that specifies the grasping target. Given the LLM’s output tends to be verbose, here we employ a special token strategy to identify those tokens that are highly relevant to the grasping target (i.e., names of target objects). Subsequently, embeddings of these identified tokens are extracted and passed through a projection network to derive a textual feature. This feature, encapsulated with essential textual information about the target, is then incorporated into the grasping prediction framework to enhance the model’s ability to accurately predict grasp poses.
However, we observe that existing grasping datasets primarily concentrate on either only visual reasoning or object grasping based on explicit instructions. To address this deficiency, we introduce a comprehensive dataset specifically designed to support the development and benchmarking of models that tackle the challenge of reasoning grasping. This dataset encompasses a diverse array of visual scenes paired with complex, implicit instructions, meticulously tailored to train and assess the reasoning and grasping capabilities of robotic systems in cluttered environments.
Our contribution can be summarized as follows:
  • We introduce a novel task of reasoning grasping, which challenges the system to not only interpret complex and implicit instructions but also to accurately generate precise grasping poses for robotic execution. This is a crucial capability for truly intelligent robotic systems.
  • We propose a benchmark model for reasoning grasping that combines multi-modal LLMs with robotic grasping techniques. Our experiments demonstrate the model’s promising performance in generating precise grasp poses in response to language instructions.
  • We have developed a unique dataset for the reasoning grasping task. It includes 64 objects, 109 parts, 1730 reasoning instructions, and around 100 million grasping poses. This dataset provides a comprehensive resource for training and evaluating models on their ability to interpret and execute complex robotic grasping tasks based on various textual instructions.
II Related Work
II-A Robotic Grasping
Robotic grasping has traditionally relied on analytical methods [29, 11, 36], which focus on understanding the geometry of objects or analyzing contact forces, and convolutional neural networks (CNNs) based methods [32, 22, 19, 24, 34, 28], which utilize large datasets of labeled grasping examples [20, 9, 14] to train models to predict grasps based on visual input. However, a critical limitation of both analytical and CNN-based methods is their lack of scene understanding and inability to process language instructions. They do not inherently know what they are grasping, which restricts their application in more dynamic, human-centric environments.
Recent advancements in robotic grasping have seen the integration of language understanding, enabling robots to grasp objects based on natural language instructions [27, 39, 41, 6, 47, 38, 46, 48, 50, 1]. This shift towards language-guided referring grasping allows robots to identify and manipulate objects as specified by users, enhancing their utility in complex scenarios. For example, Tziafas et al. focus on referring grasp synthesis, predicting grasp poses for referred objects through natural language in cluttered scenes [43]. Chen et al. proposed to learn from visual and language features jointly and predict the 2D grasp box from the RGB images [7].
Despite advancements, a critical limitation persists in robotic datasets and methods: the dependence on explicit language commands for target identification. This restricts robots’ ability to understand nuanced, implicit instructions crucial for real-world interactions. Our dataset tackles this by incorporating reasoning tasks and detailed annotations for object parts and grasping poses, enabling robots to respond to complex instructions and more accurately align with human expectations.
II-B LLMs for robotics
Recent significant breakthroughs in Large Language Models (LLMs) have been impressive [10, 5, 33, 42] in understanding and generating human-like text. The capabilities of LLMs extend well beyond text processing. Multi-modal LLMs can seamlessly integrate with other modalities, such as vision, enhancing their proficiency in tasks like visual understanding [25]. There are also a lot of efforts integrating LLMs with robotics. Many studies [2, 17, 21] have integrated LLMs into closed-loop planning structures, decomposing language-conditioned long-horizon tasks into small steps. Yet, the gap between language instructions and actions still remains. Furthermore, some studies [37, 44, 18] have employed program-like specifications to prompt LLMs, melding planning and action using a predefined library of action functions.
There is also an increase in research focused on utilizing LLMs specifically for robotic grasping. Tang et al. proposed to use LLMs to generate semantic knowledge and then selected task-oriented grasp pose from grasp candidates [40]. Mirjalili et al. employed Large Language Models to identify the optimal part of an object for grasping, treating the rest as obstacles, and then applied a traditional grasp planner to determine the best grasp [30]. In contrast to the aforementioned methods, which utilize LLMs as just one component in a multi-module grasping framework and the grasp detection relies on another grasping module, we propose a more streamlined, end-to-end approach. Our model uniquely integrates vision-language capabilities, allowing it to process language instructions and directly output the appropriate grasping pose through conversational interaction. Moreover, our model can understand implicit instructions and is capable of engaging in multi-round conversations, enhancing its applicability and intelligence in robotic grasping tasks.
III Reasoning Grasping
In this study, we introduce a novel problem of Reasoning Grasping, which involves predicting antipodal grasps for robots based on a combination of visual scenes and textual instructions. While the potential for integrating natural language into robotic manipulation is vast, its practical application in robotic grasping remains limited. Existing research [43, 46], utilizing techniques like the CLIP text encoder [35], has begun to address target-oriented grasping in cluttered environments based on straightforward language instructions, such as “Give me the banana.” However, these methods often fall short in real-world scenarios where human instructions are complex and implied, such as “I feel a little hungry.” This gap highlights the need for a more nuanced understanding and interpretation of natural language within robotic systems, driving the development of reasoning grasping.
The Reasoning Grasping task aims to generate a robotic grasp pose $g$, given an input textual instruction $t$, an input RGB image $v$, and, optionally, an input depth image $v_{d}$. The textural instruction $t$ is not limited to straightforward language instructions, it can also be complex and implicit instructions.
The output grasp poses in this work are defined in alignment with previous works [22, 32]. We represent a grasp pose as $g=\left(x,y,\theta,w,q\right)$, where $(x,y)$ represents the grasp’s central coordinates in the image plane, $\theta$ denotes the rotation angle about the $z$-axis in the range $[-\frac{\pi}{2},\frac{\pi}{2}]$, and $w$ is the grasp width of the required object in the image coordinates. The parameter $q$ signifies the grasp quality score, assigned to each point in the image on a scale from 00 to $1$, with $1$ indicating a higher likelihood of successful grasping.
IV Reasoning Grasping via Multi-modal LLMs
Figure 1: Framework of the proposed reasoning grasping model. This model processes visual images $v$ and textual instructions $t$ to output the grasp pose $g$ for the specified target object or part. The embeddings of the grasp target are passed to the grasping module for grasping poses detection.
In this section, we present an innovative model for reasoning grasping, which combines conventional robotic grasping in [22] and reasoning capabilities from multi-modal LLMs. This method effectively bridges the gap between image-based grasp prediction techniques and the nuanced understanding inherent in textual language models. Specifically, the proposed model utilizes the reasoning power of multi-modal LLMs by integrating the extracted features from its textual outputs to enhance and guide the robotic grasping prediction process.
This model addresses two key challenges: firstly, the adaptation of pre-trained multi-modal LLMs for efficient training in specific downstream tasks. While pre-trained LLMs possess strong reasoning abilities, their tendency to produce verbose and diverse outputs complicates their use in training other tasks, especially within an end-to-end framework. Secondly, the model emphasizes integrating these textual outputs from LLMs with image-based robotic grasping systems. This integration is crucial as it ensures that the rich linguistic context is effectively utilized to refine the subsequent grasping prediction.
IV-A Architecture
The proposed model consists of two main components: the reasoning module and the grasping module, as illustrated in Fig. 1. In the reasoning module, we utilize the pre-trained LLaVA [25] enhanced with LoRA [16] fine-tuning. LLaVA integrates a CLIP [35] visual encoder with LLaMA [42], an open-source Large Language Model comparable in performance to GPT-3 [5]. And LoRA, known for its computational efficiency, is incorporated into both the projection layer and all linear layers of the LLMs.
The key role of the LLMs in the reasoning module is to interpret instructions and accurately identify the target object or part for grasping. To counteract the typical verbosity of LLM outputs, we propose the use of a special token, i.e., [SPT]. This token is strategically placed around the names of the target object or the target part in the LLM’s textual outputs, i.e., [SPT] grasping target name [SPT]. The purpose of introducing this special token is to identify and isolate the crucial tokens, specifically, the names of the grasping target, from the broader textual outputs generated by LLMs.
Once the grasping target is identified within the textual outputs using the special token, we proceed to extract the last-layer embeddings of its corresponding tokens. It is important to note that there could be multiple tokens associated with the identified grasping target, and we extract the embeddings for all corresponding tokens. These extracted embeddings encapsulate the essential linguistic information related to the grasping target. Subsequently, we compute the average of these embeddings denoted as $h_{avg}$, which is then passed through a projection network to obtain the feature $f$. This feature $f$ is then fed into the grasping module, guiding the prediction of the grasp pose.
The grasping module, rooted in the CNN-based robotic grasping framework in [22], takes n-channel images as inputs. The input image is first passed through three convolutional layers and five residual layers. The extracted visual features are concatenated with the embeddings obtained from the reasoning module. Following this, we use the convolutional transpose operation to up-sample this concatenated feature, increasing its size to match the input image size. The model subsequently generates four output images, each representing different aspects of the grasp pose $g$: the grasp quality score $q$, the rotation angle expressed in $\cos 2\theta$ and $\sin 2\theta$, and the required width $w$ for the end-effector. The grasp central coordinate $(x,y)$ is determined from the output image of the grasp quality score. This is achieved by selecting the point in the image that corresponds to the maximum score.
The reasoning module processes inputs of RGB images $v$ and textual instructions $t$, while the grasping module utilizes RGB images $v$ and, optionally, depth images $v_{d}$, to accurately determine the grasp pose $g$. Specifically, the probability of generating the target grasp pose in the problem of reasoning grasping is formulated as
$p(g|v,v_{d},t)=p_{\theta_{r}}(f|v,t)\cdot p_{\theta_{g}}(g|v,v_{d},f),$ (1)
where $\theta_{r}$ and $\theta_{g}$ represent parameters in the reasoning module and the grasping module, respectively. And $f$ is the extracted feature derived from outputs of the multi-modal LLM, as illustrated in Fig. 1.
IV-B Loss Function
The proposed end-to-end model is trained by a combination of text generation loss $L_{text}$ and grasping prediction loss $L_{grasp}$. The overall objective $L$ is a weighted sum of these two components, formulated as follows:
$L=\lambda_{t}L_{text}+\lambda_{g}L_{grasp},$ (2)
where $\lambda_{t}$ and $\lambda_{g}$ are parameters that determine the relative weights of two loss terms, respectively. Here $L_{text}$ represents the auto-regressive cross-entropy loss for the textual output generated by LLMs, which assesses the performance of identifying the correct grasping target according to input instructions. On the other hand, $L_{grasp}$ evaluates the accuracy of the output grasping predictions, adhering to the loss definition detailed in [22].
IV-C Training Strategy
The proposed method leverages the pre-trained LLaVA model, with the LoRA technique for fine-tuning during the training phase. Notably, enhanced by LoRA find-tuning, the multi-modal LLM enables accurately producing textual outputs specifying the grasping target within just a few epochs. Based on this observation, our training strategy involves freezing the parameters of the LoRA-enhanced model after the initial epochs. Specifically, after the second epoch, we set the parameter $\lambda_{t}$ in the loss function to 00 and freeze the parameters of the LoRA-enhanced model. From the third epoch, our training focuses exclusively on other parameters outside the LoRA framework. This approach is designed to reduce the training time and expedite the convergence of the model, particularly in terms of grasping prediction.
Regarding the datasets used for training, our model utilizes both the reasoning grasping dataset (detailed in Section V) and the original LLaVA Instruct 150K dataset [25] utilized in LLaVA’s training. To maintain LLaVA’s visual reasoning capabilities, we initially train the model using both datasets concurrently for the first two epochs. Starting from the third epoch, with the LoRA parameters frozen, the model’s training proceeds solely with the reasoning grasping dataset.
Figure 2: Our reasoning grasping dataset enriches the GraspNet-1 billion dataset [14] with additional annotations for part-level grasping and accompanying reasoning instructions that specifically guide the grasping of corresponding objects.
Figure 3: Reasoning instruction generation. The instructions generation process involves 1) Object/Part Descriptions; 2) Automated Generation with GPT-4; and 3) Manual Review and Refinements.
V Reasoning Grasping Dataset
Reasoning grasping represents a promising yet underexplored area in robotics, offering the potential to significantly enhance how robots interact with and understand their environments. Current research and datasets primarily focus on visual reasoning or direct object manipulation based on explicit instructions, overlooking the nuanced capabilities needed for advanced understanding and response.
To bridge this gap, we introduce our reasoning grasping dataset, which builds upon the GraspNet-1 billion dataset [14]. GraspNet-1 billion consists of 190 indoor tabletop RGB-D scenes and 97,280 images captured through Kinect Azure and RealSense, featuring 88 objects. Its wide variety of scenes, objects, and grasp poses provides an excellent foundation. In our reasoning grasping dataset, we incorporate reasoning instructions alongside detailed annotations of object parts, and we exclude objects that lack semantic clarity and data associated with invalid grasps. This expanded dataset, now tailored for reasoning grasping tasks, enhances its relevance and applicability for advanced robotic manipulation research.
We provide an overview of our dataset in Fig. 2. In Section V-A, we elaborate on the process of generating reasoning instructions, a novel component of our dataset. Details of object parts annotations are discussed in Section V-B.
Figure 4: Part grasping annotation process: 1) the object’s point cloud is manually segmented into distinct parts; 2) annotated grasping poses are allocated to each segmented part of the object; and 3) the grasping pose is projected to 2D and 6D.
TABLE I: Comparison on different datasets
Grasp
Label
Modality
Multi
Object
Num.
Objects
Grasps
per Object
Num.
Grasps
Num.
Samples
Part
Annotations
Reasoning
Instructions
LISA[23] RGB N/A N/A N/A 1218
DectGPT [31] RGB N/A N/A N/A 5000
Cornell [20] Rect. RGB-D 240 33 8019 1035
Jacquard [9] Rect. RGB-D 11619 $\sim$ 20 1.1M 54,485
GraspNet [14] 6D 3D 88 $\sim$400K 1.2B 97K
OCID-Grasp [3] Rect. RGB-D 89 $\sim$ 7 75K 1763
MetaGraspNet [15] 6D 3D 82 1-5K N/A 217K
ACRONYM [13] 6D 3D 8872 2000 10.5M N/A
Grasp-Anything [45] Rect. RGB 3M 200 600M 1M
ReasoingGrasp (ours) 6D RGB-D 64 $\sim$40K $\sim$99.3M 97K
V-A Reasoning Instruction Generation
As illustrated in Fig. 3, we developed a unique and detailed method to create reasoning text instructions for our dataset by creating detailed object descriptions, automated generation with GPT-4, and meticulous manual review. The generated instructions enable robots to grasp based on nuanced, context-rich queries.
V-A1 Description Creation
The initial step in our instruction generation process involved creating detailed descriptions for each object and part, providing essential information such as functionalities, physical attributes, and typical uses. For instance, a pair of scissors was detailed as “A handheld cutting instrument with two crossing metal blades pivoted together, typically used for cutting paper or fabric. This particular pair has a black handle with yellow inner grips.” These descriptions served as the foundation for generating indirect questions and instructions.
V-A2 Automated Generation with GPT-4
The next step in instruction generation involved utilizing a customized variant of the GPT-4 model. This model was tasked with generating indirect questions and instructions for various objects and their respective parts. The specific prompt provided to the GPT was formulated to elicit responses that allude to the uses or features of an object, without directly naming it. The output consists of 10 strings: 5 indirect questions that reference the object’s functions or design, and 5 indirect instructions that describe actions involving the object or its usage in certain tasks. These are crafted to be specific, relevant, and direct yet subtly phrased, maintaining clarity and functionality.
V-A3 Manual Review and Refinement
Post-generation, each set of GPT-generated instructions underwent a rigorous manual review and refinement process. This step was crucial to ensure the instructions met our standards for subtlety, creativity, and contextual relevance. The manual intervention allowed us to fine-tune the language, enhancing the quality and applicability of the instructions for the targeted objects and their parts. This semi-automated approach balanced the efficiency of machine-generated content with the nuanced understanding and judgment of human experts.
V-B Part Annotation
V-B1 Part Segmentation
The annotation of object parts was a critical step in our dataset creation. We manually segmented the part point clouds of each object using a specially developed annotation tool. The segmentation process was guided by an understanding of how humans typically interact with and use various parts of an object. For instance, a knife was divided into its blade and handle, reflecting the distinct functional roles of each part.
V-B2 Grasping Pose Annotation
Our dataset’s annotation process includes two key components: pose assignment and pose projection, for both 3D and 2D grasping poses of each object in the scene.
Pose Assignment. Pose assignment involves aligning the grasping poses with the nearest part of the object. For each object’s segmented parts, we determine the grasping center of each point. These grasping poses are then assigned to the nearest part, ensuring an accurate association with the specific object part they correspond to.
Pose Projection. Pose projection entails transforming the grasping poses of parts into both 3D and 2D representations. This is achieved using camera matrices and object transformations. In the case of 3D, the original grasping poses are preserved in the spatial domain. For 2D projection, the 3D points of each segmented part are projected onto 2D images. This results in 2D masks that depict the spatial distribution of each part on the image plane. Grasping poses within these masks are then mapped to the corresponding parts in 2D.
These two processes of pose assignment and projection form a comprehensive approach to annotating our dataset. This approach ensures that the grasping pose data is not only detailed and extensive but also accurately aligned with the physical and functional characteristics of each object.
V-C Dataset Statistics
Our dataset significantly enhances the GraspNet-1B foundation by annotating object parts and incorporating reasoning instructions for robotic grasping tasks and detailed object parts annotations.
In ensuring data quality, low-quality grasping poses were eliminated based on confidence evaluations and objects lacking semantic information were removed. In total, our enhanced dataset comprises an extensive collection of 1,730 instruction pairs, 64 objects, 109 segmented parts, and around 100 million associated grasping poses.
Table I shows the statistics comparison of the previous grasping and visual reasoing datasets with ours. Our dataset distinguishes itself from existing visual reasoning works by integrating actionable robotic manipulation data, a feature not commonly found in similar datasets. Moreover, unlike other grasping datasets, we offer detailed reasoning instructions and part-specific grasping information. This unique combination of features positions our dataset as a critical resource for developing more intelligent and adaptable robotic systems, bridging the gap between visual reasoning and practical robotic tasks.
VI Experiment
In this section, we report the experiment results of the proposed method on the reasoning grasping dataset and real-world experiments. In the following experiments, we utilize both explicit instructions, such as “Pick up the <object>”, as well as implicit instructions that do not directly mention the grasping target.
VI-A Implementation Details
For the multi-modal LLM in the proposed model, we utilize the pre-trained LLaVA-7B-v0, which is derived from the large language model LLaMA-7B. As for the LoRA fine-tuning, we choose a rank $r=64$. During the training, we use the batch size of $8$, and the learning rate is set to $45e-$ 4, utilizing a cosine annealing schedule for adaptive learning rate adjustment. For the relative weight parameters in the loss function 2, we set $\lambda_{t}=1$ and $\lambda_{g}=1$ for the initial configuration, and starting from the third epoch, the parameter $\lambda_{t}$ is adjusted to 00.
VI-B Baselines
For the problem of reasoning grasping, we implement two baselines to compare with our proposed model. One baseline employs a CLIP text encoder [35] to extract textual features, while the other utilizes a modular approach incorporating the latest LLaVA model.
VI-B1 CLIP $+$ GR-ConvNet Baseline
In this baseline, we combine a CLIP text encoder with a CNN-based grasp detection model, GR-ConvNet [22]. We first utilize the CLIP text encoder to extract the textual feature from input instructions, and this feature is then integrated into the hidden layers of GR-ConvNet. While CLIP can effectively extract language features relevant to visual content, it may fall short in capturing implicit reasoning compared to the VLM.
VI-B2 LLaVa $\rightarrow$ GR-ConvNet Baseline
This is a modular baseline that integrates the latest LLaVA-1.6-34b [26] with a pre-trained grasping detection model GR-ConvNet. As LLaVA cannot directly output the grasping pose, this baseline operates in two stages to obtain the grasping pose: first, LLaVA takes an image and an instruction as the input and is prompted to output a bounding box of the target object in the image. Then GR-ConvNet detects the optimal grasping pose within this specified bounding box. Note that LLaVA-1.6-34b has a much higher performance than LLaVa-v0-7b [25], which is the base model of our proposed reasoning grasping model.
VI-C Results
VI-C1 Text Generation
We first evaluate the precision of text generation by the fine-tuned multi-modal LLM in the reasoning module. Accurately generating special tokens and target names is crucial, since this identified grasping target will be utilized in the subsequent grasping module. The special tokens and target names act as intermediate states and enable the model to eventually generate a precise grasping pose from given instructions. We differentiate between explicit instructions, which directly mention the object’s name, and implicit instructions, which hint at the target object without stating its name. Note that the instructions used for testing were not seen during the training phase. Results, presented in Table II, show that our model can successfully interpret $92.52\%$ and $88.79\%$ of implicit instructions for object grasping and part grasping, respectively. This demonstrates our model’s good reasoning capabilities in understanding the grasping instructions.
TABLE II: Text generation accuracy.
Explicit Instruction Implicit Instruction
Object Grasping 99.01% 92.52%
Part Grasping 95.33% 88.79%
TABLE III: Grasp prediction accuracy on the reasoning grasping dataset.
Explicit Instruction Implicit Instruction Explicit Instruction Implicit Instruction
Object Grasping Object Grasping Part Grasping Part Grasping
Models
R@1
R@2
R@3
R@1
R@2
R@3
R@1
R@2
R@3
R@1
R@2
R@3
LLaVA $\rightarrow$ GR-ConvNet
26.80%
39.17%
44.66%
18.34%
23.76%
31.29%
15.38%
23.52%
37.93%
0.92%
3.84%
8.53%
CLIP $+$ GR-ConvNet
50.40%
62.88%
66.38%
44.35%
57.15%
61.09%
43.78%
55.71%
59.61%
28.88%
42.34%
47.37%
Reasoning Grasping 64.49% 84.11% 86.92% 63.55% 73.83% 77.57% 59.81% 70.01% 81.31% 61.68% 71.96% 83.18%
LLaVA $\rightarrow$ GR-ConvNet (with depth)
31.88%
40.72%
45.95%
20.39%
25.77%
34.84%
20.11%
29.30%
36.63%
1.36%
2.77%
11.95%
CLIP $+$ GR-ConvNet (with depth)
54.11%
66.50%
69.57%
45.89%
58.11%
62.50%
45.53%
59.51%
62.27%
33.06%
48.55%
52.25%
Reasoning Grasping (with depth) 60.75% 71.95% 76.63% 57.94% 70.09% 77.46% 69.16% 72.90% 77.55% 64.48% 75.70% 80.37%
VI-C2 Reasoning Grasping Result
Table III shows the grasp prediction results on the reasoning grasping dataset. Performance is reported using the rectangle evaluation metric [19]. A grasp pose is deemed valid if it fulfills the following two conditions: 1) The Intersection over Union (IoU) score between the predicted and ground truth rectangles exceeds $25\%$; 2) The angular deviation between the orientations of the predicted and ground truth rectangles is less than 30 degrees. We reported R@$k$, indicating the presence of valid grasps within the top-$k$ grasp predictions. Four distinct scenarios are evaluated, categorized by the combination of instruction types (explicit or implicit) and grasping targets (object or part). We assess both baseline models and our reasoning grasping model with and without depth input. Note that when depth information is utilized in the reasoning grasping model, it only serves as input for the grasping module and is not incorporated into the reasoning module, as shown in Fig. 1.
The results show that our reasoning grasping model outperforms the LLaVA $\rightarrow$ GR-ConvNet and CLIP $+$ GR-ConvNet baselines across all scenarios. The first baseline, LLaVA $\rightarrow$ GR-ConvNet, shows limited effectiveness, as the direct combining of two models fails to convey sufficient information for accurate grasp detection. Although the CLIP $+$ GR-ConvNet baseline exhibits some improvements, it still falls short, particularly in scenarios involving part grasping with implicit instructions. This indicates CLIP’s limited capability in processing implicit instructions. In contrast, our model excels in all scenarios, both with and without depth input, showcasing its superior ability to interpret implicit instructions and accurately generate the corresponding grasping poses. Note that some common failure cases are distinguishing objects with similar appearances, such as a shampoo bottle and a hair conditioner bottle, or identifying specific parts of complex objects like the uniformly colored toy camel.
VI-C3 Visual Reasoning Ability
Besides reasoning grasping, we also examine the visual reasoning capabilities of our model. The objective is to determine whether the model, after being fine-tuned, maintains the visual comprehension skills of the original LLaVA model. For a quantitative evaluation of performance, we employ GPT-4 to assess the quality of responses generated by our model. This is inspired by the evaluation methods in [25, 8]. For each given image and question pair, both the original LLaVa model and our fine-tuned version provide answers related to the input image. We subsequently submit the questions, and answers from both LLaVA and our model, along with the ground-truth text descriptions, to a text-only GPT-4-based judge. This judge assesses the responses based on their helpfulness, relevance, accuracy, and level of detail in comparison to the ground-truth descriptions. GPT-4 then allocates an overall performance score ranging from 0 to 10, where higher scores denote superior performance.
The image-question-answer sets used in this evaluation are randomly chosen from the LLaVa-Instruct-150k dataset [25]. According to the judge of GPT-4, the original LLaVA-7B-v0 achieves $8.25/10$ when taking ground-truth textual descriptions as references. On the other hand, our model achieves $7.43/10$ under the same GPT-4 judge. This shows that our fine-tuned model still preserves a good portion of the reasoning ability of the original LLaVA.
VI-D Real World Experiments
Figure 5: Real world experiments with implicit instructions.
We conduct real-world experiments to evaluate the proposed reasoning grasping model and the CLIP $+$ GR-ConvNet baseline. A 7-DoF Franka Emika Panda equipped with a Franka Hand parallel gripper was utilized for grasping execution. To perceive the objects, one Azure Kinect was positioned in front of the robot. A cropped $480\times 480$ RGB image of the workspace was utilized as the input to the model. To execute the 2D grasps generated from the model, we always apply a top-down grasp with the height computed using the depth of the grasping point. In our experiments, we randomly place 4-8 objects on a tabletop and provide either explicit or implicit language instruction to the robot to grasp an object or a specific part, as shown in Fig. 5. We select objects that are either identical or similar to those in the dataset. We deliberately excluded any objects or parts that are not infeasible for grasping, such as items too wide for the robot’s gripper. The performance is evaluated using three evaluation metrics: the correctness of generating special tokens and grasping target names, the accuracy of the output grasp pose, and the success in lifting the object. The results are reported in Table IV.
Overall our proposed model can generate accurate special tokens and target names 39 out of 40 trials. This again shows the excellent reasoning capabilities of our model to interpret the various instructions. The failure case involved the model incorrectly identifying a screwdriver instead of scissors in response to a prompt requesting a tool to open a package box. In terms of grasping performance, we assessed top-3 output grasp poses and manually selected the most feasible one for execution on the robot arm. The CLIP $+$ GR-ConvNet baseline and our reasoning grasping model produced 15/40 and 25/40 accurate grasping poses, respectively, with the success rates for lifting objects are 9/40 and 16/40, respectively. The experiment results demonstrate that our model outperforms the baseline in four distinct scenarios, showing its superior ability in reasoning grasping tasks.
TABLE IV: Real-world experiment results.
CLIP + GR-ConvNet Reasoning Grasping
Object Grasping with Explicit Instruction
Token Accuracy
N/A 10/10
Pose Accuracy
5/10 7/10
Execution Success
2/10 5/10
Object Grasping with Implicit Instruction
Token Accuracy
N/A 9/10
Pose Accuracy
3/10 5/10
Execution Success
2/10 3/10
Part Grasping with Explicit Instruction
Token Accuracy
N/A 10/10
Pose Accuracy
5/10 7/10
Execution Success
3/10 4/10
Part Grasping with Implicit Instruction
Token Accuracy
N/A 10/10
Pose Accuracy
2/10 6/10
Execution Success
2/10 4/10
VII Conclusion
In this paper, we introduce a novel task of reasoning grasping, where robots need to interpret and generate grasping poses based on implicit instructions. By leveraging the reasoning ability in a multi-modal large language model, our proposed model can understand indirect verbal commands and generate corresponding grasping poses. In addition, we also present the first dataset for reasoning grasping, derived from the GraspNet-1 billion dataset, featuring implicit human instructions alongside object and part grasping annotations. Experiment results demonstrate that our approach can effectively comprehend implicit instructions and accurately generate corresponding grasping poses. Overall, this work offers valuable insights bridging implicit human instructions and generating precise robot actions.
While the proposed reasoning grasping model shows promising results, the model is limited when generating optimal grasping poses for novel objects and refining poses through conversational interactions. A possible reason is that the variety of objects with grasping annotations is still limited. For future work, the LLaVA-v0-7b model is not the state-of-the-art Visual Language Model. Other models such as LLaVA-1.6 [26] or GPT-4V [49] have shown much better performance and stronger ability in terms of visual understanding. Our model’s architecture, with its modular reasoning and grasping components, allows for flexibility in upgrading to more sophisticated models in the future.
Appendix A Object and Part Statistics
TABLE V: Selected Objects and Parts for the Training Split of the Dataset
Object Name Part Name
6D Grasps
per Sample
Rect Grasps
per Sample
Num.
Scenes
0 Cracker Box - 10971 4672 35
1 Tomato Soup Can - 16625 8284 34
5 Banana Stem, Flesh 1298 2941 25
7 Red Mug Handle, Rim, Body 3363 2045 25
8 Power Drill Handle, Body, Batterypack, Chuck 2445 4476 26
9 Scissor Handle, Blade 548 574 22
11 Strawberry - 130 903 24
14 Peach - 960 1922 25
15 Pear - 948 2069 26
17 Orange - 581 1466 32
18 Knife Handle, Blade 1975 1526 29
20 Red Screwdriver Handle, Shaft 1191 1910 29
21 Racquetball - 443 573 6
22 Blue Cup Rim, Body 3960 2770 29
36 Daobao Wash Soap Cap, Body 2594 1744 25
37 Nzskincare Mouth Rinse Cap, Body 1806 1786 24
38 Daobao Sod Cap, Body 3256 2227 25
40 Kispa Cleanser Cap, Body 931 1654 25
41 Darlie Toothpaste Cap, Body 442 1181 25
43 Baoke Marker Cap, Body 953 697 21
44 Hosjam Toothpaste Pump Pump, Body 725 1958 22
46 Dish Rim, Body 8 22 25
48 Camel Legs, Head, Body 96 663 25
51 Elephant Legs, Head, Body, Truck 218 1692 24
52 Rhinocero Legs, Head, Body, Horn 152 1419 22
57 Weiquan Chicken Bouillon Cap, Body 2297 2617 25
58 Darlie Toothpaste Box - 5526 1646 25
60 Black Mouse - 170 618 25
61 Dabao Facewash Cap, Body 4683 2297 22
62 Pantene Shampoo Cap, Body 502 1175 26
63 Head Shoulders Supreme Cap, Body 777 1700 24
66 Head Shoulders Care Cap, Body 1338 1269 24
70 Tape - 4465 1859 25
Total 33 54 76377 64355 100
Appendix B 3D Object and Part Visualization
Figure 6: 3D Visualization of Selected Objects and its Parts
Appendix C Prompt for Instruction Generation
Figure 7: Example Prompts used for Initial Reasoning Instruction Generation with GPT-4

### v2

@@
 The utilization of CNN-based neural networks has demonstrated efficiency in generating high-quality grasping poses from visual input [31, 21, 18, 23, 33, 27].
-One significant limitation of these methods is the lack of scene understanding, resulting in robots being unable to identify the objects they are grasping.
-Recent developments have introduced methods that interpret language instructions such as “I want a stapler” [7] or “Pick the food box in front of the ball” [43] for targeted grasping, enabling robots to pick objects as directed.
+One significant limitation of these methods is the lack of scene understanding.
+The robot cannot identify the objects they are grasping.
+Recent methods have introduced language-guided grasping with instructions such as “I want a stapler” [6] or “Pick the food box in front of the ball” [42].
 However, those methods will struggle when dealing with implicit instructions, where the desired object is not explicitly named.
-For example, if a user says “I need to drink water”, an intelligent assistant robot should autonomously identify and grasp a cup from a selection of household items on the table, although “cup” does not appear in the user’s instruction.
-Such scenarios, prevalent in real-world setting, underscore the importance of advanced understanding and autonomy in robotic grasping — an area that remains a significant challenge.
-Recently, the advance of large language models like ChatGPT [33] has introduced new possibilities for language reasoning, significantly enhancing our ability to process and interpret complex language structures.
-Beyond text, numerous studies have leveraged multi-modal Large Language Models (LLMs) for enhanced visual understanding and interaction [51, 12, 25], enabling these models to interpret and generate responses based on a combination of textual and visual inputs.
-However, these approaches have been confined largely to visual comprehension and text generation, lacking the capability to generate precise robotic actions.
-While there have been efforts to integrate LLMs with robotics [2, 17, 21, 37, 44, 18], these have predominantly focused on high-level planning or have modularized the LLM for reasoning while relying on traditional methods for downstream actions.
-Brohan et al.
-[4] utilize a Vision Language Model (VLM) to transfer web knowledge to robot control, but it requires extensive efforts to train on a large amount of data.
-In this work, we introduce the novel task of Reasoning Grasping, where the robot is tasked with generating appropriate grasping poses based on implicit instructions.
-This task presents two significant challenges, primarily due to the nuanced requirements of interpreting complex, context-rich instructions and accurately generating the corresponding grasp poses.
-To address these challenges, we propose a model that is uniquely designed to directly output grasp poses according to visual scenes and complex instructions in cluttered environments.
-Specifically, this model initially leverages a multi-modal LLM to interpret both images and instructions, producing a textual output that specifies the grasping target.
-Given the LLM’s output tends to be verbose, here we employ a special token strategy to identify those tokens that are highly relevant to the grasping target (i.e., names of target objects).
-Subsequently, embeddings of these identified tokens are extracted and passed through a projection network to derive a textual feature.
-This feature, encapsulated with essential textual information about the target, is then incorporated into the grasping prediction framework to enhance the model’s ability to accurately predict grasp poses.
-However, we observe that existing grasping datasets primarily concentrate on either only visual reasoning or object grasping based on explicit instructions.
-To address this deficiency, we introduce a comprehensive dataset specifically designed to support the development and benchmarking of models that tackle the challenge of reasoning grasping.
-This dataset encompasses a diverse array of visual scenes paired with complex, implicit instructions, meticulously tailored to train and assess the reasoning and grasping capabilities of robotic systems in cluttered environments.
+For example, if a user says “I need to drink water”, an intelligent assistant robot should identify and grasp a cup from a selection of household items, although “cup” does not appear in the user’s instruction.
+Such scenarios are common in real-world settings and the solutions to them remain open questions.
+Recently, the advance of large language models like ChatGPT [32] has introduced new possibilities for language reasoning.
+Beyond text, numerous studies have leveraged multi-modal Large Language Models (LLMs) for visual understanding [50, 11, 24], enabling these models to interpret and generate responses based on both textual and visual inputs.
+However, these approaches focus on visual comprehension and text generation, lacking the capability to generate robotic actions.
+While there have been efforts to integrate LLMs with robotics [2, 16, 20, 36, 43, 17], these concentrate on high-level planning.
+In this work, we introduce the novel task of Reasoning Grasping, where a robot determines grasp poses based on users’ “implicit” instructions.
+“Implicit” instructions are language instructions that do not specify the name of the grasp target.
+A formal definition of reasoning grasping and implicit instruction is given in Sec III.
+To address the reasoning grasping task, we propose a model that directly output grasp poses according to image inputs and language instructions in cluttered environments.
+Specifically, this model leverages a multi-modal LLM to interpret both images and instructions.
+We employ a special token strategy to identify tokens that are relevant to the grasping target (i.e., names of target objects).
+Subsequently, embeddings of these identified tokens are used to generate accurate grasping poses.
+To train this model, we extend the GraspNet-1 billion dataset [13] with diverse implicit instructions and object part grasping.
 Our contribution can be summarized as follows:
-  • We introduce a novel task of reasoning grasping, which challenges the system to not only interpret complex and implicit instructions but also to accurately generate precise grasping poses for robotic execution.
-This is a crucial capability for truly intelligent robotic systems.
-  • We propose a benchmark model for reasoning grasping that combines multi-modal LLMs with robotic grasping techniques.
-Our experiments demonstrate the model’s promising performance in generating precise grasp poses in response to language instructions.
+  • We introduce a novel task of reasoning grasping, which directs robots to grasp objects based on implicit instructions.
+  • We propose a multi-modal LLM for reasoning grasping. Our experiments demonstrate the model’s promising performance.
   • We have developed a unique dataset for the reasoning grasping task.
 It includes 64 objects, 109 parts, 1730 reasoning instructions, and around 100 million grasping poses.
-This dataset provides a comprehensive resource for training and evaluating models on their ability to interpret and execute complex robotic grasping tasks based on various textual instructions.
+This dataset provides a comprehensive resource for training and evaluating models on reasoning grasping tasks.
 II Related Work
 II-A Robotic Grasping
-Robotic grasping has traditionally relied on analytical methods [29, 11, 36], which focus on understanding the geometry of objects or analyzing contact forces, and convolutional neural networks (CNNs) based methods [32, 22, 19, 24, 34, 28], which utilize large datasets of labeled grasping examples [20, 9, 14] to train models to predict grasps based on visual input.
+Robotic grasping has traditionally relied on analytical methods [28, 10, 35], which focus on understanding the geometry of objects or analyzing contact forces.
+Convolutional neural networks (CNNs) based methods [31, 21, 18, 23, 33, 27] utilize large datasets of labeled grasping examples [19, 8, 13] to train models to predict grasps based on visual input.
 However, a critical limitation of both analytical and CNN-based methods is their lack of scene understanding and inability to process language instructions.
@@
 Recent advancements in robotic grasping have seen the integration of language understanding, enabling robots to grasp objects based on natural language instructions [26, 38, 40, 5, 46, 37, 45, 47, 49, 1].
-This shift towards language-guided referring grasping allows robots to identify and manipulate objects as specified by users, enhancing their utility in complex scenarios.
+This shift allows robots to identify and manipulate objects as specified by users, enhancing their utility in complex scenarios.
 For example, Tziafas et al.
@@
 proposed to learn from visual and language features jointly and predict the 2D grasp box from the RGB images [6].
-Despite advancements, a critical limitation persists in robotic datasets and methods: the dependence on explicit language commands for target identification.
-This restricts robots’ ability to understand nuanced, implicit instructions crucial for real-world interactions.
-Our dataset tackles this by incorporating reasoning tasks and detailed annotations for object parts and grasping poses, enabling robots to respond to complex instructions and more accurately align with human expectations.
 II-B LLMs for robotics
-Recent significant breakthroughs in Large Language Models (LLMs) have been impressive [10, 5, 33, 42] in understanding and generating human-like text.
-The capabilities of LLMs extend well beyond text processing.
+Recent breakthroughs in Large Language Models (LLMs) have been impressive [9, 4, 32, 41] in understanding and generating human-like text.
 Multi-modal LLMs can seamlessly integrate with other modalities, such as vision, enhancing their proficiency in tasks like visual understanding [24].
@@
 Tang et al.
-proposed to use LLMs to generate semantic knowledge and then selected task-oriented grasp pose from grasp candidates [40].
+proposed GraspGPT to use LLMs to generate semantic knowledge and then selected task-oriented grasp pose from grasp candidates [39].
 Mirjalili et al.
-employed Large Language Models to identify the optimal part of an object for grasping, treating the rest as obstacles, and then applied a traditional grasp planner to determine the best grasp [30].
-In contrast to the aforementioned methods, which utilize LLMs as just one component in a multi-module grasping framework and the grasp detection relies on another grasping module, we propose a more streamlined, end-to-end approach.
-Our model uniquely integrates vision-language capabilities, allowing it to process language instructions and directly output the appropriate grasping pose through conversational interaction.
-Moreover, our model can understand implicit instructions and is capable of engaging in multi-round conversations, enhancing its applicability and intelligence in robotic grasping tasks.
+proposed LAN-grasp to identify the optimal part of an object for grasping with LLM, treating the rest as obstacles, and then applied a traditional grasp planner to determine the best grasp [29].
+Compared to GraspGPT and LAN-grasp, our method is distinguished by its ability to operate in cluttered scenes and its end-to-end training framework, while GraspGPT and LAN-grasp utilize modular frameworks and depend heavily on other pre-trained grasp detection models.
 III Reasoning Grasping
-In this study, we introduce a novel problem of Reasoning Grasping, which involves predicting antipodal grasps for robots based on a combination of visual scenes and textual instructions.
-While the potential for integrating natural language into robotic manipulation is vast, its practical application in robotic grasping remains limited.
-Existing research [43, 46], utilizing techniques like the CLIP text encoder [35], has begun to address target-oriented grasping in cluttered environments based on straightforward language instructions, such as “Give me the banana.” However, these methods often fall short in real-world scenarios where human instructions are complex and implied, such as “I feel a little hungry.” This gap highlights the need for a more nuanced understanding and interpretation of natural language within robotic systems, driving the development of reasoning grasping.
-The Reasoning Grasping task aims to generate a robotic grasp pose $g$, given an input textual instruction $t$, an input RGB image $v$, and, optionally, an input depth image $v_{d}$.
-The textural instruction $t$ is not limited to straightforward language instructions, it can also be complex and implicit instructions.
+In this section, we define the task of Reasoning Grasping, which aims to generate a robotics grasp pose $g$, given an input textual instruction $t$, an input RGB image $v$, and, optionally, an input depth image $v_{d}$.
+In the reasoning grasping task, instructions $t$ are not limited to being straightforward, they can also be “implicit”.
+We define implicit instructions as language instructions that do not explicitly specify the name of the grasp target but provide relevant contextual cues or indicate users’ intent implying the grasp target.
+For example, implicit instructions may include cues such as shapes, colors, or other attributes of grasp targets, or they may describe users’ needs or intent implying the grasp targets based on their functionalities.
 The output grasp poses in this work are defined in alignment with previous works [21, 31].
@@
 The parameter $q$ signifies the grasp quality score, assigned to each point in the image on a scale from 00 to $1$, with $1$ indicating a higher likelihood of successful grasping.
-IV Reasoning Grasping via Multi-modal LLMs
 Figure 1: Framework of the proposed reasoning grasping model.
@@
 The embeddings of the grasp target are passed to the grasping module for grasping poses detection.
+IV Reasoning Grasping via Multi-modal LLMs
 In this section, we present an innovative model for reasoning grasping, which combines conventional robotic grasping in [21] and reasoning capabilities from multi-modal LLMs.
-This method effectively bridges the gap between image-based grasp prediction techniques and the nuanced understanding inherent in textual language models.
-Specifically, the proposed model utilizes the reasoning power of multi-modal LLMs by integrating the extracted features from its textual outputs to enhance and guide the robotic grasping prediction process.
-This model addresses two key challenges: firstly, the adaptation of pre-trained multi-modal LLMs for efficient training in specific downstream tasks.
-While pre-trained LLMs possess strong reasoning abilities, their tendency to produce verbose and diverse outputs complicates their use in training other tasks, especially within an end-to-end framework.
-Secondly, the model emphasizes integrating these textual outputs from LLMs with image-based robotic grasping systems.
+Specifically, the proposed model utilizes the reasoning power of multi-modal LLMs by integrating the extracted features from its textual outputs into the robotic grasping prediction process.
+This model addresses two key challenges: firstly, it addresses the adaptation of pre-trained multi-modal LLMs for efficient training in downstream tasks.
+Although LLMs have strong reasoning abilities, they often produce verbose and diverse outputs, making it challenging to directly integrate them into training other tasks, especially within an end-to-end framework.
+Secondly, the model addresses the combination of textual outputs from LLMs and image-based robotic grasping systems.
 This integration is crucial as it ensures that the rich linguistic context is effectively utilized to refine the subsequent grasping prediction.
 IV-A Architecture
-The proposed model consists of two main components: the reasoning module and the grasping module, as illustrated in Fig.
+The proposed model framework is illustrated in Fig.
 1.
-In the reasoning module, we utilize the pre-trained LLaVA [25] enhanced with LoRA [16] fine-tuning.
+As for multi-modal LLMs, we utilize the pre-trained LLaVA [24] enhanced with LoRA [15] fine-tuning techniques.
 LLaVA integrates a CLIP [34] visual encoder with LLaMA [41], an open-source Large Language Model comparable in performance to GPT-3 [4].
-And LoRA, known for its computational efficiency, is incorporated into both the projection layer and all linear layers of the LLMs.
-The key role of the LLMs in the reasoning module is to interpret instructions and accurately identify the target object or part for grasping.
-To counteract the typical verbosity of LLM outputs, we propose the use of a special token, i.e., [SPT].
-This token is strategically placed around the names of the target object or the target part in the LLM’s textual outputs, i.e., [SPT] grasping target name [SPT].
-The purpose of introducing this special token is to identify and isolate the crucial tokens, specifically, the names of the grasping target, from the broader textual outputs generated by LLMs.
-Once the grasping target is identified within the textual outputs using the special token, we proceed to extract the last-layer embeddings of its corresponding tokens.
-It is important to note that there could be multiple tokens associated with the identified grasping target, and we extract the embeddings for all corresponding tokens.
+And LoRA, known for its computational efficiency, is incorporated into both the projection layer and all linear layers of multi-modal LLMs.
+The key role of multi-modal LLMs in the proposed model is to interpret both instruction $t$ and image $v$, then accurately identify the grasping target.
+Such a target could be an object or, more challenging, a specific part within cluttered scenes.
+To address the verbosity of LLM outputs, we introduce a special token, i.e., [SPT], strategically placed around the names of the grasping target in the LLM textual outputs.
+For instance, [SPT] grasping target name [SPT].
+The purpose of introducing this special token is to identify and isolate the crucial tokens, specifically, the names of the grasping target, from the textual outputs generated by LLMs.
+Once the grasping target is identified using the special token, we proceed to extract the last-layer embeddings of its corresponding tokens.
+It is important to note that there could be multiple tokens associated with the identified grasping target, and we extract embeddings for all identified tokens.
 These extracted embeddings encapsulate the essential linguistic information related to the grasping target.
-Subsequently, we compute the average of these embeddings denoted as $h_{avg}$, which is then passed through a projection network to obtain the feature $f$.
-This feature $f$ is then fed into the grasping module, guiding the prediction of the grasp pose.
-The grasping module, rooted in the CNN-based robotic grasping framework in [22], takes n-channel images as inputs.
+We then compute the average of these embeddings denoted as $h_{avg}$, which is passed through a projection network to obtain the feature $f$.
+This feature $f$ is then fed into the image-based grasping detection process, guiding the prediction of the grasp pose.
+The grasping detection in our model, rooted in the CNN-based robotic grasping framework in [21], takes n-channel images as inputs.
 The input image is first passed through three convolutional layers and five residual layers.
@@
 The grasp central coordinate $(x,y)$ is determined from the output image of the grasp quality score.
-This is achieved by selecting the point in the image that corresponds to the maximum score.
-The reasoning module processes inputs of RGB images $v$ and textual instructions $t$, while the grasping module utilizes RGB images $v$ and, optionally, depth images $v_{d}$, to accurately determine the grasp pose $g$.
-Specifically, the probability of generating the target grasp pose in the problem of reasoning grasping is formulated as
-$p(g|v,v_{d},t)=p_{\theta_{r}}(f|v,t)\cdot p_{\theta_{g}}(g|v,v_{d},f),$ (1)
-where $\theta_{r}$ and $\theta_{g}$ represent parameters in the reasoning module and the grasping module, respectively.
-And $f$ is the extracted feature derived from outputs of the multi-modal LLM, as illustrated in Fig.
-1.
+This is obtained by selecting the point in the image with the highest quality score.
+Notably, the depth image $v_{d}$ is an optional input modality for the grasping prediction, depending on data availability.
 IV-B Loss Function
@@
 The proposed method leverages the pre-trained LLaVA model, with the LoRA technique for fine-tuning during the training phase.
-Notably, enhanced by LoRA find-tuning, the multi-modal LLM enables accurately producing textual outputs specifying the grasping target within just a few epochs.
+With LoRA fine-tuning, the multi-modal LLM can generate the grasp target within just a few epochs.
 Based on this observation, our training strategy involves freezing the parameters of the LoRA-enhanced model after the initial epochs.
-Specifically, after the second epoch, we set the parameter $\lambda_{t}$ in the loss function to 00 and freeze the parameters of the LoRA-enhanced model.
-From the third epoch, our training focuses exclusively on other parameters outside the LoRA framework.
+Specifically, from the third epoch, we set the parameter $\lambda_{t}$ in the loss function to 00 and freeze the parameters of the LoRA-enhanced model.
+Our training then focuses exclusively on other parameters outside the LoRA framework.
 This approach is designed to reduce the training time and expedite the convergence of the model, particularly in terms of grasping prediction.
@@
 Starting from the third epoch, with the LoRA parameters frozen, the model’s training proceeds solely with the reasoning grasping dataset.
-Figure 2: Our reasoning grasping dataset enriches the GraspNet-1 billion dataset [14] with additional annotations for part-level grasping and accompanying reasoning instructions that specifically guide the grasping of corresponding objects.
+Figure 2: Our reasoning grasping dataset enriches the GraspNet-1 billion dataset [13] with additional annotations for part-level grasping and reasoning instructions.
 Figure 3: Reasoning instruction generation. The instructions generation process involves 1) Object/Part Descriptions; 2) Automated Generation with GPT-4; and 3) Manual Review and Refinements.
 V Reasoning Grasping Dataset
-Reasoning grasping represents a promising yet underexplored area in robotics, offering the potential to significantly enhance how robots interact with and understand their environments.
-Current research and datasets primarily focus on visual reasoning or direct object manipulation based on explicit instructions, overlooking the nuanced capabilities needed for advanced understanding and response.
-To bridge this gap, we introduce our reasoning grasping dataset, which builds upon the GraspNet-1 billion dataset [14].
-GraspNet-1 billion consists of 190 indoor tabletop RGB-D scenes and 97,280 images captured through Kinect Azure and RealSense, featuring 88 objects.
+To train the model for reasoning grasping tasks, we introduce our reasoning grasping dataset, which builds upon the GraspNet-1 billion dataset [13].
+GraspNet-1 billion consists of 190 indoor tabletop RGB-D scenes, 88 objects, and 97280 images.
 Its wide variety of scenes, objects, and grasp poses provides an excellent foundation.
-In our reasoning grasping dataset, we incorporate reasoning instructions alongside detailed annotations of object parts, and we exclude objects that lack semantic clarity and data associated with invalid grasps.
-This expanded dataset, now tailored for reasoning grasping tasks, enhances its relevance and applicability for advanced robotic manipulation research.
-We provide an overview of our dataset in Fig.
+We extend the GraspNet-1B dataset with reasoning instructions as well as object parts grasping annotations.
+We provide some examples in our dataset in Fig.
 2.
-In Section V-A, we elaborate on the process of generating reasoning instructions, a novel component of our dataset.
-Details of object parts annotations are discussed in Section V-B.
 Figure 4: Part grasping annotation process: 1) the object’s point cloud is manually segmented into distinct parts; 2) annotated grasping poses are allocated to each segmented part of the object; and 3) the grasping pose is projected to 2D and 6D.
@@
 As illustrated in Fig.
-3, we developed a unique and detailed method to create reasoning text instructions for our dataset by creating detailed object descriptions, automated generation with GPT-4, and meticulous manual review.
+3, we developed a unique and detailed method to create reasoning text instructions for our dataset by creating detailed object descriptions, automated generation with GPT-4, and manual review.
 The generated instructions enable robots to grasp based on nuanced, context-rich queries.
@@
 The next step in instruction generation involved utilizing a customized variant of the GPT-4 model.
-This model was tasked with generating indirect questions and instructions for various objects and their respective parts.
-The specific prompt provided to the GPT was formulated to elicit responses that allude to the uses or features of an object, without directly naming it.
+This model was tasked with generating indirect questions and instructions for various objects and their parts.
 The output consists of 10 strings: 5 indirect questions that reference the object’s functions or design, and 5 indirect instructions that describe actions involving the object or its usage in certain tasks.
-These are crafted to be specific, relevant, and direct yet subtly phrased, maintaining clarity and functionality.
+These are crafted to be specific, relevant, and direct for grasping tasks.
 V-A3 Manual Review and Refinement
-Post-generation, each set of GPT-generated instructions underwent a rigorous manual review and refinement process.
-This step was crucial to ensure the instructions met our standards for subtlety, creativity, and contextual relevance.
+Post-generation, each set of GPT-generated instructions underwent a manual review and refinement process.
 The manual intervention allowed us to fine-tune the language, enhancing the quality and applicability of the instructions for the targeted objects and their parts.
-This semi-automated approach balanced the efficiency of machine-generated content with the nuanced understanding and judgment of human experts.
+This semi-automated approach balanced the efficiency of machine-generated content and judgment of human experts.
 V-B Part Annotation
 V-B1 Part Segmentation
-The annotation of object parts was a critical step in our dataset creation.
 We manually segmented the part point clouds of each object using a specially developed annotation tool.
@@
 V-B2 Grasping Pose Annotation
-Our dataset’s annotation process includes two key components: pose assignment and pose projection, for both 3D and 2D grasping poses of each object in the scene.
-Pose Assignment.
+Our dataset’s annotation process includes two key components: pose assignment and pose projection, for both 3D and 2D grasping poses, as shown in Fig. 4.
 Pose assignment involves aligning the grasping poses with the nearest part of the object.
@@
 These grasping poses are then assigned to the nearest part, ensuring an accurate association with the specific object part they correspond to.
-Pose Projection.
-Pose projection entails transforming the grasping poses of parts into both 3D and 2D representations.
+Pose projection transforms the grasping poses of parts into both 3D and 2D representations.
 This is achieved using camera matrices and object transformations.
@@
 Grasping poses within these masks are then mapped to the corresponding parts in 2D.
-These two processes of pose assignment and projection form a comprehensive approach to annotating our dataset.
-This approach ensures that the grasping pose data is not only detailed and extensive but also accurately aligned with the physical and functional characteristics of each object.
 V-C Dataset Statistics
-Our dataset significantly enhances the GraspNet-1B foundation by annotating object parts and incorporating reasoning instructions for robotic grasping tasks and detailed object parts annotations.
 In ensuring data quality, low-quality grasping poses were eliminated based on confidence evaluations and objects lacking semantic information were removed.
 In total, our enhanced dataset comprises an extensive collection of 1,730 instruction pairs, 64 objects, 109 segmented parts, and around 100 million associated grasping poses.
-Table I shows the statistics comparison of the previous grasping and visual reasoing datasets with ours.
-Our dataset distinguishes itself from existing visual reasoning works by integrating actionable robotic manipulation data, a feature not commonly found in similar datasets.
-Moreover, unlike other grasping datasets, we offer detailed reasoning instructions and part-specific grasping information.
-This unique combination of features positions our dataset as a critical resource for developing more intelligent and adaptable robotic systems, bridging the gap between visual reasoning and practical robotic tasks.
+Table I shows the statistics comparison of the previous grasping and visual reasoning datasets with ours.
+Unlike other grasping datasets, we offer detailed reasoning instructions and part-specific grasping information.
 VI Experiment
@@
 During the training, we use the batch size of $8$, and the learning rate is set to $45e-$ 4, utilizing a cosine annealing schedule for adaptive learning rate adjustment.
-For the relative weight parameters in the loss function 2, we set $\lambda_{t}=1$ and $\lambda_{g}=1$ for the initial configuration, and starting from the third epoch, the parameter $\lambda_{t}$ is adjusted to 00.
+For the relative weight parameters in the loss function 1, we set $\lambda_{t}=1$ and $\lambda_{g}=1$ for the initial configuration, and starting from the third epoch, the parameter $\lambda_{t}$ is adjusted to 00.
+For the train-test-split, we use $90\%$ of the first 100 scenes in GraspNet-1 billion [13] for training and $10\%$ of the first 100 scenes for testing.
+We sampled 50k from the original LLaVA Instruct 150K dataset [24] to mix with 100k grasping data to maintain the visual reasoning capabilities.
 VI-B Baselines
@@
 Note that LLaVA-1.6-34b has a much higher performance than LLaVa-v0-7b [24], which is the base model of our proposed reasoning grasping model.
+This baseline mirrors a similar process to LAN-grasp.
 VI-C Results
@@
 Object Grasping Object Grasping Part Grasping Part Grasping
-Models
-R@1
-R@2
-R@3
-R@1
-R@2
-R@3
-R@1
-R@2
-R@3
-R@1
-R@2
-R@3
-LLaVA $\rightarrow$ GR-ConvNet
-26.80%
-39.17%
-44.66%
-18.34%
-23.76%
-31.29%
-15.38%
-23.52%
-37.93%
-0.92%
-3.84%
-8.53%
-CLIP $+$ GR-ConvNet
-50.40%
-62.88%
-66.38%
-44.35%
-57.15%
-61.09%
-43.78%
-55.71%
-59.61%
-28.88%
-42.34%
-47.37%
+Models R@1 R@2 R@3 R@1 R@2 R@3 R@1 R@2 R@3 R@1 R@2 R@3
+LLaVA $\rightarrow$ GR-ConvNet 26.80% 39.17% 44.66% 18.34% 23.76% 31.29% 15.38% 23.52% 37.93% 0.92% 3.84% 8.53%
+CLIP $+$ GR-ConvNet 50.40% 62.88% 66.38% 44.35% 57.15% 61.09% 43.78% 55.71% 59.61% 28.88% 42.34% 47.37%
 Reasoning Grasping 64.49% 84.11% 86.92% 63.55% 73.83% 77.57% 59.81% 70.01% 81.31% 61.68% 71.96% 83.18%
-LLaVA $\rightarrow$ GR-ConvNet (with depth)
-31.88%
-40.72%
-45.95%
-20.39%
-25.77%
-34.84%
-20.11%
-29.30%
-36.63%
-1.36%
-2.77%
-11.95%
-CLIP $+$ GR-ConvNet (with depth)
-54.11%
-66.50%
-69.57%
-45.89%
-58.11%
-62.50%
-45.53%
-59.51%
-62.27%
-33.06%
-48.55%
-52.25%
+LLaVA $\rightarrow$ GR-ConvNet (with depth) 31.88% 40.72% 45.95% 20.39% 25.77% 34.84% 20.11% 29.30% 36.63% 1.36% 2.77% 11.95%
+CLIP $+$ GR-ConvNet (with depth) 54.11% 66.50% 69.57% 45.89% 58.11% 62.50% 45.53% 59.51% 62.27% 33.06% 48.55% 52.25%
 Reasoning Grasping (with depth) 60.75% 71.95% 76.63% 57.94% 70.09% 77.46% 69.16% 72.90% 77.55% 64.48% 75.70% 80.37%
@@
 This shows that our fine-tuned model still preserves a good portion of the reasoning ability of the original LLaVA.
+It also shows that the multi-round conversation ability is preserved, as the image-question-answer sets include sequences of multi-round conversations.
 VI-D Real World Experiments
@@
 The failure case involved the model incorrectly identifying a screwdriver instead of scissors in response to a prompt requesting a tool to open a package box.
+We classified it as a failure solely because it did not match the specific pairing in our dataset, which serves as the basis for our evaluation metrics.
+This example underscores that even when the model’s response is technically incorrect, it is still reasonable.
 In terms of grasping performance, we assessed top-3 output grasp poses and manually selected the most feasible one for execution on the robot arm.
+We conducted 10 trials for each scenario, totaling 80 trials, to compare reasoning grasping with CLIP $+$ GR-ConvNet baseline.
 The CLIP $+$ GR-ConvNet baseline and our reasoning grasping model produced 15/40 and 25/40 accurate grasping poses, respectively, with the success rates for lifting objects are 9/40 and 16/40, respectively.
+The main reasons for accurate poses but fail to lift were due to the object slipping from the gripper or the object being too heavy.
 The experiment results demonstrate that our model outperforms the baseline in four distinct scenarios, showing its superior ability in reasoning grasping tasks.

### v3

@@
 In this study, we introduce a novel task: reasoning grasping, where robots need to generate grasp poses based on indirect verbal instructions or intentions.
-To accomplish this, we propose an end-to-end reasoning grasping model that integrates a multi-modal Large Language Model (LLM) with a vision-based robotic grasping framework.
-In addition, we present the first reasoning grasping benchmark dataset generated from the GraspNet-1 billion, incorporating implicit instructions for object-level and part-level grasping, and this dataset will soon be available for public access.
+To accomplish this, we propose an end-to-end reasoning grasping model that integrates a multimodal Large Language Model (LLM) with a vision-based robotic grasping framework.
+In addition, we present the first reasoning grasping benchmark dataset generated from the GraspNet-1 billion, incorporating implicit instructions for object-level and part-level grasping.
 Our results show that directly integrating CLIP or LLaVA with the grasp detection model performs poorly on the challenging reasoning grasping tasks, while our proposed model demonstrates significantly enhanced performance both in the reasoning grasping benchmark and real-world experiments.
-I Introduction
+11footnotetext: Work done while the authors were with Baidu Research.
+Keywords: Robotics Grasping, Multimodal Large Language Model
+1 Introduction
 Robotic grasping has long been a subject of extensive study.
@@
 While there have been efforts to integrate LLMs with robotics [13, 14, 15, 16, 17, 18], these concentrate on high-level planning.
-In this work, we introduce the novel task of Reasoning Grasping, where a robot determines grasp poses based on users’ “implicit” instructions.
+Figure 1: Overview.
+We integrate the reasoning abilities of multi-modal Large Language Models with robotic grasping.
+The resulting model interprets complex and implicit instructions, accurately predicts robotic grasping poses for target objects or specific parts within cluttered environments, and supports multi-round conversations with users.
+In the textual output from the model, the grasping target is indicated by two special tokens [SPT], as demonstrated in the figure.
+The output grasp poses are visualized in the images with rectangles.
+In this work, we introduce the novel task of Reasoning Grasping (Fig.
+1), where a robot determines grasp poses based on users’ “implicit” instructions.
 “Implicit” instructions are language instructions that do not specify the name of the grasp target.
-A formal definition of reasoning grasping and implicit instruction is given in Sec III.
+A formal definition of reasoning grasping and implicit instruction is given in Sec 3.
 To address the reasoning grasping task, we propose a model that directly output grasp poses according to image inputs and language instructions in cluttered environments.
@@
 To train this model, we extend the GraspNet-1 billion dataset [19] with diverse implicit instructions and object part grasping.
-Our contribution can be summarized as follows:
-  • We introduce a novel task of reasoning grasping, which directs robots to grasp objects based on implicit instructions.
-  • We propose a multi-modal LLM for reasoning grasping. Our experiments demonstrate the model’s promising performance.
-  • We have developed a unique dataset for the reasoning grasping task.
-It includes 64 objects, 109 parts, 1730 reasoning instructions, and around 100 million grasping poses.
+Our contribution can be summarized as follows: (1) We introduce a novel task of reasoning grasping, which directs robots to grasp objects based on implicit instructions.
+(2) We propose a multi-modal LLM for reasoning grasping.
+Our experiments demonstrate the model’s promising performance.
+(3) We have developed a unique dataset for the reasoning grasping task.
+It includes 64 objects, 109 parts, 1,730 reasoning instructions, and around 100 million grasping poses.
 This dataset provides a comprehensive resource for training and evaluating models on reasoning grasping tasks.
-II Related Work
-II-A Robotic Grasping
+2 Related Work
+Robotic Grasping.
 Robotic grasping has traditionally relied on analytical methods [20, 21, 22], which focus on understanding the geometry of objects or analyzing contact forces.
 Convolutional neural networks (CNNs) based methods [1, 2, 3, 4, 5, 6] utilize large datasets of labeled grasping examples [23, 24, 19] to train models to predict grasps based on visual input.
-However, a critical limitation of both analytical and CNN-based methods is their lack of scene understanding and inability to process language instructions.
+A critical limitation of both analytical and CNN-based methods is their lack of scene understanding and inability to process language instructions.
 They do not inherently know what they are grasping, which restricts their application in more dynamic, human-centric environments.
-Recent advancements in robotic grasping have seen the integration of language understanding, enabling robots to grasp objects based on natural language instructions [26, 38, 40, 5, 46, 37, 45, 47, 49, 1].
+Recent advancements in robotic grasping have seen the integration of language understanding, enabling robots to grasp objects based on natural language instructions [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
 This shift allows robots to identify and manipulate objects as specified by users, enhancing their utility in complex scenarios.
-For example, Tziafas et al.
-focus on referring grasp synthesis, predicting grasp poses for referred objects through natural language in cluttered scenes [42].
-Chen et al.
-proposed to learn from visual and language features jointly and predict the 2D grasp box from the RGB images [6].
-II-B LLMs for robotics
-Recent breakthroughs in Large Language Models (LLMs) have been impressive [9, 4, 32, 41] in understanding and generating human-like text.
+However, those models require explicit object names as the input and struggle to understand complex and implicit language instructions.
+LLMs for Robotics.
+Large Language Models (LLMs) [36, 37, 9, 38] are impressive in understanding and generating human-like text.
 Multi-modal LLMs can seamlessly integrate with other modalities, such as vision, enhancing their proficiency in tasks like visual understanding [12].
 There are also a lot of efforts integrating LLMs with robotics.
-Many studies [2, 16, 20] have integrated LLMs into closed-loop planning structures, decomposing language-conditioned long-horizon tasks into small steps.
+Many studies [13, 14, 15, 39] have integrated LLMs into closed-loop planning structures, decomposing language-conditioned long-horizon tasks into small steps.
 Yet, the gap between language instructions and actions still remains.
-Furthermore, some studies [36, 43, 17] have employed program-like specifications to prompt LLMs, melding planning and action using a predefined library of action functions.
-There is also an increase in research focused on utilizing LLMs specifically for robotic grasping.
+Furthermore, some studies [16, 17, 18, 40] have employed program-like specifications to prompt LLMs, melding planning and action using a predefined library of action functions.
+There is also an increase in research focused on utilizing LLMs for robotic grasping.
 Tang et al.
@@
 proposed LAN-grasp to identify the optimal part of an object for grasping with LLM, treating the rest as obstacles, and then applied a traditional grasp planner to determine the best grasp [42].
-Compared to GraspGPT and LAN-grasp, our method is distinguished by its ability to operate in cluttered scenes and its end-to-end training framework, while GraspGPT and LAN-grasp utilize modular frameworks and depend heavily on other pre-trained grasp detection models.
-III Reasoning Grasping
+Compared to GraspGPT and LAN-grasp, which utilize modular frameworks that depend heavily on other pre-trained grasp detection models, our method utilizes an end-to-end training framework and can operate in cluttered scenes.
+3 Reasoning Grasping
 In this section, we define the task of Reasoning Grasping, which aims to generate a robotics grasp pose $g$, given an input textual instruction $t$, an input RGB image $v$, and, optionally, an input depth image $v_{d}$.
@@
 For example, implicit instructions may include cues such as shapes, colors, or other attributes of grasp targets, or they may describe users’ needs or intent implying the grasp targets based on their functionalities.
-The output grasp poses in this work are defined in alignment with previous works [21, 31].
-We represent a grasp pose as $g=\left(x,y,\theta,w,q\right)$, where $(x,y)$ represents the grasp’s central coordinates in the image plane, $\theta$ denotes the rotation angle about the $z$-axis in the range $[-\frac{\pi}{2},\frac{\pi}{2}]$, and $w$ is the grasp width of the required object in the image coordinates.
-The parameter $q$ signifies the grasp quality score, assigned to each point in the image on a scale from 00 to $1$, with $1$ indicating a higher likelihood of successful grasping.
+Some application scenarios are “Users cannot visually confirm the robot’s surroundings” or “Tasks that require internal knowledge for decision making”.
+We include the motivations and detailed applications scenarios in Appendix A.
+4 Reasoning Grasping via Multi-modal LLMs
 Figure 2: Framework of the proposed reasoning grasping model.
@@
 The embeddings of the grasp target are passed to the grasping module for grasping poses detection.
-IV Reasoning Grasping via Multi-modal LLMs
-In this section, we present an innovative model for reasoning grasping, which combines conventional robotic grasping in [21] and reasoning capabilities from multi-modal LLMs.
-Specifically, the proposed model utilizes the reasoning power of multi-modal LLMs by integrating the extracted features from its textual outputs into the robotic grasping prediction process.
-This model addresses two key challenges: firstly, it addresses the adaptation of pre-trained multi-modal LLMs for efficient training in downstream tasks.
-Although LLMs have strong reasoning abilities, they often produce verbose and diverse outputs, making it challenging to directly integrate them into training other tasks, especially within an end-to-end framework.
-Secondly, the model addresses the combination of textual outputs from LLMs and image-based robotic grasping systems.
-This integration is crucial as it ensures that the rich linguistic context is effectively utilized to refine the subsequent grasping prediction.
-IV-A Architecture
+In this section, we present our model for reasoning grasping that utilizes the reasoning ability of multi-modal LLMs by integrating the extracted features from its textual outputs into the robotic grasping prediction process.
+4.1 Architecture
 The proposed model framework is illustrated in Fig.
-1.
+2.
 As for multi-modal LLMs, we utilize the pre-trained LLaVA [12] enhanced with LoRA [43] fine-tuning techniques.
@@
 Following this, we use the convolutional transpose operation to up-sample this concatenated feature, increasing its size to match the input image size.
-The model subsequently generates four output images, each representing different aspects of the grasp pose $g$: the grasp quality score $q$, the rotation angle expressed in $\cos 2\theta$ and $\sin 2\theta$, and the required width $w$ for the end-effector.
+The model subsequently generates four output images, each representing different aspects of the grasp pose $g=\left(x,y,\theta,w,q\right)$: the grasp quality score $q$, the rotation angle expressed in $\cos 2\theta$ and $\sin 2\theta$, and the required width $w$ for the end-effector.
+The definition of the grasp pose aligns with previous work [2, 1].
 The grasp central coordinate $(x,y)$ is determined from the output image of the grasp quality score.
@@
 Notably, the depth image $v_{d}$ is an optional input modality for the grasping prediction, depending on data availability.
-IV-B Loss Function
+4.2 Loss Function
 The proposed end-to-end model is trained by a combination of text generation loss $L_{text}$ and grasping prediction loss $L_{grasp}$.
@@
 On the other hand, $L_{grasp}$ evaluates the accuracy of the output grasping predictions, adhering to the loss definition detailed in [2].
-IV-C Training Strategy
+4.3 Training Strategy
 The proposed method leverages the pre-trained LLaVA model, with the LoRA technique for fine-tuning during the training phase.
@@
 This approach is designed to reduce the training time and expedite the convergence of the model, particularly in terms of grasping prediction.
-Regarding the datasets used for training, our model utilizes both the reasoning grasping dataset (detailed in Section V) and the original LLaVA Instruct 150K dataset [24] utilized in LLaVA’s training.
+Regarding the datasets used for training, our model utilizes both the reasoning grasping dataset (detailed in Section 5) and the original LLaVA Instruct 150K dataset [12] utilized in LLaVA’s training.
 To maintain LLaVA’s visual reasoning capabilities, we initially train the model using both datasets concurrently for the first two epochs.
 Starting from the third epoch, with the LoRA parameters frozen, the model’s training proceeds solely with the reasoning grasping dataset.
-Figure 2: Our reasoning grasping dataset enriches the GraspNet-1 billion dataset [13] with additional annotations for part-level grasping and reasoning instructions.
-Figure 3: Reasoning instruction generation. The instructions generation process involves 1) Object/Part Descriptions; 2) Automated Generation with GPT-4; and 3) Manual Review and Refinements.
-V Reasoning Grasping Dataset
+5 Reasoning Grasping Dataset
 To train the model for reasoning grasping tasks, we introduce our reasoning grasping dataset, which builds upon the GraspNet-1 billion dataset [19].
-GraspNet-1 billion consists of 190 indoor tabletop RGB-D scenes, 88 objects, and 97280 images.
+GraspNet-1 billion consists of 190 indoor tabletop RGB-D scenes, 88 objects, and 97,280 images.
 Its wide variety of scenes, objects, and grasp poses provides an excellent foundation.
 We extend the GraspNet-1B dataset with reasoning instructions as well as object parts grasping annotations.
-We provide some examples in our dataset in Fig.
+We provide some examples in our dataset in Appendix F.
+5.1 Reasoning Instruction Generation
+We developed a unique and detailed method to create reasoning text instructions for our dataset by creating detailed object descriptions, automated generation with GPT-4, and manual review (Appendix D).
+The generated instructions enable robots to grasp based on nuanced, context-rich queries.
+Description Creation.
+The initial step in our instruction generation process involved creating detailed descriptions for each object and part, providing essential information such as functionalities, physical attributes, and typical uses.
+For instance, a pair of scissors was detailed as “A handheld cutting instrument with two crossing metal blades pivoted together, typically used for cutting paper or fabric.
+This particular pair has a black handle with yellow inner grips.” These descriptions served as the foundation for generating indirect questions and instructions.
+Automated Generation with GPT-4.
+The next step involved generating instructions automatically using GPT-4.
+An example prompt is provided in Appendix C.
+GPT-4 was employed to generate indirect questions and instructions for various objects and their parts.
+The output consists of 10 strings: 5 indirect questions that reference the object’s functions or design, and 5 indirect instructions that describe actions involving the object or its usage in certain tasks.
+These are crafted to be specific, relevant, and direct for grasping tasks.
+Manual Review and Refinement.
+Post-generation, each set of GPT-generated instructions underwent a manual review and refinement process.
+The manual intervention allowed us to fine-tune the language, enhancing the quality and applicability of the instructions for the targeted objects and their parts.
+This semi-automated approach balanced the efficiency of machine-generated content and judgment of human experts.
+5.2 Part Annotation
+We manually segmented the part point clouds of each object using a specially developed annotation tool.
+The segmentation process was guided by an understanding of how humans typically interact with and use various parts of an object.
+For instance, a knife was divided into its blade and handle, reflecting the distinct functional roles of each part.
+The ground truth grasping poses are assigned to the nearest part.
+We include the details of part annotation in Appendix E.
+5.3 Dataset Statistics
+In ensuring data quality, low-quality grasping poses were eliminated based on confidence evaluations and objects lacking semantic information were removed.
+In total, our enhanced dataset comprises an extensive collection of 1,730 instruction pairs, 64 objects, 109 segmented parts, and around 100 million associated grasping poses.
+Unlike other grasping datasets, we offer detailed reasoning instructions and part-specific grasping information.
+6 Experiments
+In this section, we report the experiment results of the proposed method on the reasoning grasping dataset and real-world experiments.
+In the following experiments, we utilize both explicit instructions, such as “Pick up the <object>”, as well as implicit instructions that do not directly mention the grasping target.
+6.1 Baselines
+For the problem of reasoning grasping, we implement two baselines to compare with our proposed model.
+One baseline employs a CLIP text encoder [44] to extract textual features, while the other utilizes a modular approach incorporating the latest LLaVA model.
+CLIP $+$ GR-ConvNet Baseline.
+In this baseline, we combine a CLIP text encoder with a CNN-based grasp detection model, GR-ConvNet [2].
+We first utilize the CLIP text encoder to extract the textual feature from input instructions, and this feature is then integrated into the hidden layers of GR-ConvNet.
+While CLIP can effectively extract language features relevant to visual content, it may fall short in capturing implicit reasoning compared to the VLM.
+LLaVa $\rightarrow$ GR-ConvNet Baseline.
+This is a modular baseline that integrates the latest LLaVA-1.6-34b [45] with a pre-trained grasping detection model GR-ConvNet.
+As LLaVA cannot directly output the grasping pose, this baseline operates in two stages to obtain the grasping pose: first, LLaVA takes an image and an instruction as the input and is prompted to output a bounding box of the target object in the image.
+Then GR-ConvNet detects the optimal grasping pose within this specified bounding box.
+Note that LLaVA-1.6-34b has a much higher performance than LLaVa-v0-7b [12], which is the base model of our proposed reasoning grasping model.
+This baseline mirrors a similar process to LAN-grasp.
+6.2 Results
+Table 1: Text generation accuracy.
+Explicit Instruction Implicit Instruction
+Object Grasping 99.01% 92.52%
+Part Grasping 95.33% 88.79%
+Text Generation.
+We first evaluate the precision of text generation by the fine-tuned multi-modal LLM in the reasoning module.
+Accurately generating special tokens and target names is crucial, since this identified grasping target will be utilized in the subsequent grasping module.
+We differentiate between explicit instructions, which directly mention the object’s name, and implicit instructions, which hint at the target object without stating its name.
+Note that the instructions used for testing were not seen during the training phase.
+Results, presented in Table 1, show that our model can successfully interpret $92.52\%$ and $88.79\%$ of implicit instructions for object grasping and part grasping, respectively.
+This demonstrates our model’s good reasoning capabilities in understanding the grasping instructions.
+Table 2: Grasp prediction accuracy on the reasoning grasping dataset.
+Explicit Instruction Implicit Instruction Explicit Instruction Implicit Instruction
+Object Grasping Object Grasping Part Grasping Part Grasping
+Models R@1 R@3 R@1 R@3 R@1 R@3 R@1 R@3
+LLaVA $\rightarrow$ GR-ConvNet 26.80% 44.66% 18.34% 31.29% 15.38% 37.93% 0.92% 8.53%
+CLIP $+$ GR-ConvNet 50.40% 66.38% 44.35% 61.09% 43.78% 59.61% 28.88% 47.37%
+Reasoning Grasping 64.49% 86.92% 63.55% 77.57% 59.81% 81.31% 61.68% 83.18%
+LLaVA $\rightarrow$ GR-ConvNet (with depth) 31.88% 45.95% 20.39% 34.84% 20.11% 36.63% 1.36% 11.95%
+CLIP $+$ GR-ConvNet (with depth) 54.11% 69.57% 45.89% 62.50% 45.53% 62.27% 33.06% 52.25%
+Reasoning Grasping (with depth) 60.75% 76.63% 57.94% 77.46% 69.16% 77.55% 64.48% 80.37%
+Reasoning Grasping Result.
+Table 2 shows the grasp prediction results on the reasoning grasping dataset.
+Performance is reported using the rectangle evaluation metric [3].
+A grasp pose is deemed valid if it fulfills the following two conditions: 1) The Intersection over Union (IoU) score between the predicted and ground truth rectangles exceeds $25\%$; 2) The angular deviation between the orientations of the predicted and ground truth rectangles is less than 30 degrees.
+We reported R@$k$, indicating the presence of valid grasps within the top-$k$ grasp predictions.
+Four distinct scenarios are evaluated, categorized by the combination of instruction types (explicit or implicit) and grasping targets (object or part).
+We assess both baseline models and our reasoning grasping model with and without depth input.
+Note that when depth information is utilized in the reasoning grasping model, it only serves as input for the grasping module and is not incorporated into the reasoning module, as shown in Fig.
 2.
+The results show that our reasoning grasping model outperforms the LLaVA $\rightarrow$ GR-ConvNet and CLIP $+$ GR-ConvNet baselines across all scenarios.
+The first baseline, LLaVA $\rightarrow$ GR-ConvNet, shows limited effectiveness, as the direct combining of two models fails to convey sufficient information for accurate grasp detection.
+Although the CLIP $+$ GR-ConvNet baseline exhibits some improvements, it still falls short, particularly in scenarios involving part grasping with implicit instructions.
+This indicates CLIP’s limited capability in processing implicit instructions.
+In contrast, our model excels in all scenarios, both with and without depth input, showcasing its superior ability to interpret implicit instructions and accurately generate the corresponding grasping poses.
+Note that some common failure cases are distinguishing objects with similar appearances, such as a shampoo bottle and a hair conditioner bottle, or identifying specific parts of complex objects like the uniformly colored toy camel.
+Visual Reasoning Ability.
+Besides reasoning grasping, we also examine the visual reasoning capabilities of our model.
+The objective is to determine whether the model, after being fine-tuned, maintains the visual comprehension skills of the original LLaVA model.
+For a quantitative evaluation of performance, we employ a GPT-4 judge to assess the quality of responses generated by our model.
+This is inspired by the evaluation methods in [12, 46].
+According to the judge of GPT-4, the original LLaVA-7B-v0 achieves $8.25/10$ when taking ground-truth textual descriptions as references.
+On the other hand, our model achieves $7.43/10$ under the same GPT-4 judge.
+This shows that our fine-tuned model still preserves a good portion of the reasoning ability of the original LLaVA.
+We include the details of the experiments in Appendix J.
+6.3 Real World Experiments
+Figure 3: Real world experiments with implicit instructions.
+We conduct real-world experiments to evaluate the proposed reasoning grasping model and the CLIP $+$ GR-ConvNet baseline.
+A 7-DoF Franka Emika Panda equipped with a Franka Hand parallel gripper was utilized for grasping execution.
+To perceive the objects, one Azure Kinect was positioned in front of the robot.
+A cropped $480\times 480$ RGB image of the workspace was utilized as the input to the model.
+To execute the 2D grasps generated from the model, we always apply a top-down grasp with the height computed using the depth of the grasping point.
+In our experiments, we randomly place 4-8 objects on a tabletop and provide either explicit or implicit language instruction to the robot to grasp an object or a specific part, as shown in Fig.
+3.
+We select objects that are either identical or similar to those in the dataset.
+We deliberately excluded any objects or parts that are not infeasible for grasping, such as items too wide for the robot’s gripper.
+The performance is evaluated using three evaluation metrics: the correctness of generating special tokens and grasping target names, the accuracy of the output grasp pose, and the success in lifting the object.
+The results are reported in Table 3.
+Table 3: Real-world experiment results.
+Model Token Accuracy Pose Accuracy Execution Success
+Object / Explicit CLIP + GR-ConvNet N/A 5/10 2/10
+Reasoning Grasping 10/10 7/10 5/10
+Object / Implicit CLIP + GR-ConvNet N/A 3/10 2/10
+Reasoning Grasping 9/10 5/10 3/10
+Part / Explicit CLIP + GR-ConvNet N/A 5/10 3/10
+Reasoning Grasping 10/10 7/10 4/10
+Part / Implicit CLIP + GR-ConvNet N/A 2/10 2/10
+Reasoning Grasping 10/10 6/10 4/10
+Overall our proposed model can generate accurate special tokens and target names 39 out of 40 trials.
+This again shows the excellent reasoning capabilities of our model to interpret the various instructions.
+The failure case involved the model incorrectly identifying a screwdriver instead of scissors in response to a prompt requesting a tool to open a package box.
+We classified it as a failure solely because it did not match the specific pairing in our dataset, which serves as the basis for our evaluation metrics.
+This example underscores that even when the model’s response is technically incorrect, it is still reasonable.
+In terms of grasping performance, we assessed top-3 output grasp poses and manually selected the most feasible one for execution on the robot arm.
+We conducted 10 trials for each scenario, totaling 80 trials, to compare reasoning grasping with CLIP $+$ GR-ConvNet baseline.
+The CLIP $+$ GR-ConvNet baseline and our reasoning grasping model produced 15/40 and 25/40 accurate grasping poses, respectively, with the success rates for lifting objects are 9/40 and 16/40, respectively.
+The main reasons for accurate poses but fail to lift were due to the object slipping from the gripper or the object being too heavy.
+The experiment results demonstrate that our model outperforms the baseline in four distinct scenarios, showing its superior ability in reasoning grasping tasks.
+It’s important to emphasize the unique challenges posed by our study, which involve “implicit” instructions, grasping in cluttered environments, and task-oriented objects/parts grasping.
+This level of complexity has not been studied before.
+As a comparison, a recent study [8] deals with grasping in cluttered environments with “explicit” instructions achieves $60\%$ for grounding accuracy (accurate poses) and $20\%$ for success rate (success lift), respectively.
+7 Conclusion
+In this paper, we introduce a novel task of reasoning grasping, where robots need to interpret and generate grasping poses based on implicit instructions.
+By leveraging the reasoning ability in a multi-modal large language model, our proposed model can understand indirect verbal commands and generate corresponding grasping poses.
+In addition, we also present the first dataset for reasoning grasping, derived from the GraspNet-1 billion dataset, featuring implicit human instructions alongside object and part grasping annotations.
+Experiment results demonstrate that our approach can effectively comprehend implicit instructions and accurately generate corresponding grasping poses.
+Overall, this work offers valuable insights bridging implicit human instructions and generating precise robot actions.
+While the proposed reasoning grasping model shows promising results, the model is limited when generating optimal grasping poses for novel objects and refining poses through conversational interactions.
+A possible reason is that the variety of objects with grasping annotations is still limited.
+For future work, the LLaVA-v0-7b model is not the state-of-the-art Visual Language Model.
+Other models such as LLaVA-1.6 [45] or GPT-4V [47] have shown much better performance and stronger ability in terms of visual understanding.
+Our model’s architecture, with its modular reasoning and grasping components, allows for flexibility in upgrading to more sophisticated models in the future.
+Acknowledgments
+We would like to express our gratitude to the reviewers for their insightful feedback and constructive comments.
+Additionally, we thank Zhixian Ye for his assistance with the experiments during the paper rebuttal.
+Appendix A Motivations and Application Scenarios
+We aim to advance robotic systems to not only understand direct human commands but to also possess the reasoning capability to interpret implicit instructions.
+These implicit instructions can provide a general description of users’ needs, requiring the robot to infer the grasping target on its own, without explicitly naming the object.
+This is very common in real-world scenarios.
+Here are several practical application scenarios where direct instructions may be unavailable:
+  • Users cannot visually confirm the robot’s surroundings: For example, blind people can instruct the robot to “find a sweater in a warm color that matches the pants”.
+The robot could also be asked to “find something to hold these papers together”, where it might choose from a stapler, paper clips, or binder clips based on availability.
+  • Tasks that require internal knowledge for decision making: Robots can utilize their internal knowledge for tasks like sorting in recycling facilities.
+They could be instructed to “sort materials that are recyclable and compostable”, identifying items based on texture and material type, which facilitates sorting without the need to know whether the objects are recyclable or not.
+  • Naming objects is impractical or complex: In some situations, explicitly naming objects is impractical or the names are complex and difficult to remember.
+For instance, telling a robot, “I need my morning medicine”, allows the robot to use its routine knowledge to fetch the correct medication without the user needing to recall specific drug names, which are usually long and complex.
+  • Managing multiple items: Explicitly naming multiple objects can be cumbersome and inefficient.
+In emergency medical situations, a medical robot might be instructed to “gather all items necessary for suturing wounds”.
+It would need to discern which tools and supplies are relevant, such as needles, thread, and antiseptics, based on the medical context provided.
+For general tidying tasks in home organization, a home robot could be instructed with phrases like “The living room is messy, tidy up by picking up toys”, allowing the robot to identify and collect toys without naming each item individually (e.g., teddy bear, toy train, etc.).
+Appendix B Compared to Other Methods.
+Here we provide detailed comparisons with similar approaches: GraspGPT [41], LAN-grasp [42], and GPT-4v [48].
+Figure 4: Comparison of our method with GraspGPT and LAN-grasp.
+B.1 Compared to GraspGPT and LAN-grasp.
+Clutted scenes.
+GraspGPT and LAN-grasp are designed to grasp specific parts of single objects, whereas our method focuses on grasping in cluttered scenes.
+During our experiments, we intentionally tested our model in cluttered scenes to assess its ability to reason about grasping targets.
+These targets could be entire objects or, more challenging, specific parts within cluttered scenes.
+End-to-end trained.
+One significant distinction is that our model is an end-to-end system for directly generating numerical grasp predictions.
+However, GraspGPT and LAN-grasp have modular frameworks that both employ external pre-trained models for grasp detection.
+  • GraspGPT, illustrated in Figure 4, has to take grasp candidates as the input.
+GraspGPT serves as a grasp evaluator, selecting a grasp pose from given input grasp candidates, which are generated from pre-trained models.
+This reliance on frozen models designed for single objects limits GraspGPT’s adaptability to complex environments, such as cluttered scenes.
+  • LAN-grasp generates the bounding boxes of targets using Vision-Language Models (e.g., OWL-Vit [49]) and then uses GraspIt!
+[50] for determining the output grasp pose.
+This process without any trainable model heavily depends on the performance of utilized pre-trained models.
+Also, it is challenging to adapt LAN-grasp to new tasks or scenes.
+Moreover, our baseline“LLaVA $\rightarrow$ GR-ConvNet” mirrors a similar process to LAN-grasp, utilizing LLaVA [12] for bounding box identification and a pre-trained GR-ConvNet [2] for grasp pose detection.
+However, its performance is unsatisfactory when applied to cluttered scenes on our reasoning-grasping dataset, as shown in Table III of the paper.
+In contrast, our model jointly trains language models and grasping detection models, eliminating the need for external pre-trained models.
+Experimental results (Table III) demonstrate that our end-to-end approach outperforms modular frameworks such as “LLaVA $\rightarrow$ GR-ConvNet”.
+B.2 Compared to GPT-4v.
+While models like GPT-4 with vision (GPT-4v) [48] show potential in tasks such as object detection, they still face challenges in new tasks with generating unique numerical predictions like grasp poses, despite careful prompting.
+To illustrate GPT-4v’s ability to generate numerical predictions for novel tasks like robotic grasping, we conduct additional experiments using both zero-shot and few-shot prompting techniques.
+Specifically, we use images from the benchmark Cornell Grasp dataset [4], where each image contains only a single object.
+GPT-4v is prompted to generate grasp poses for the object in the scene.
+As shown in Figure 5, adapting GPT-4v to new grasp detection tasks, even in simple scenarios with single objects, remains challenging despite careful prompt design and providing examples.
+Figure 5: Grasping detection with few-shot prompting using GPT-4v on Cornell Grasp dataset.
+Appendix C Prompt for Instruction Generation
+Figure 6: Example Prompts used for Initial Reasoning Instruction Generation with GPT-4
+Appendix D Reasoning Instruction Generation
+Figure 7: Reasoning instruction generation. The instructions generation process involves 1) Object/Part Descriptions; 2) Automated Generation with GPT-4; and 3) Manual Review and Refinements.
+Appendix E Part Segmentation and Grasping Pose Annotation
+Our dataset’s annotation process includes two key components: pose assignment and pose projection, for both 3D and 2D grasping poses, as shown in Fig.
+8.
+Pose assignment involves aligning the grasping poses with the nearest part of the object.
+For each object’s segmented parts, we determine the grasping center of each point.
+These grasping poses are then assigned to the nearest part, ensuring an accurate association with the specific object part they correspond to.
+Pose projection transforms the grasping poses of parts into both 3D and 2D representations.
+This is achieved using camera matrices and object transformations.
+In the case of 3D, the original grasping poses are preserved in the spatial domain.
+For 2D projection, the 3D points of each segmented part are projected onto 2D images.
+This results in 2D masks that depict the spatial distribution of each part on the image plane.
+Grasping poses within these masks are then mapped to the corresponding parts in 2D.
 Figure 8: Part grasping annotation process: 1) the object’s point cloud is manually segmented into distinct parts; 2) annotated grasping poses are allocated to each segmented part of the object; and 3) the grasping pose is projected to 2D and 6D.
-TABLE I: Comparison on different datasets
+Appendix F Examples of Reasoning Grasping Dataset
+Figure 9: Our reasoning grasping dataset enriches the GraspNet-1 billion dataset [19] with additional annotations for part-level grasping and reasoning instructions.
+Appendix G Comparison with Different Datasets
+Table 4: Comparison with different datasets
 Grasp
@@
 Instructions
-LISA[22] RGB N/A N/A N/A 1218
-DectGPT [30] RGB N/A N/A N/A 5000
-Cornell [19] Rect. RGB-D 240 33 8019 1035
-Jacquard [8] Rect. RGB-D 11619 $\sim$ 20 1.1M 54,485
+LISA[51] RGB N/A N/A N/A 1,218
+DectGPT [52] RGB N/A N/A N/A 5,000
+Cornell [23] Rect. RGB-D 240 33 8,019 1,035
+Jacquard [24] Rect. RGB-D 11,619 $\sim$ 20 1.1M 54,485
 GraspNet [19] 6D 3D 88 $\sim$400K 1.2B 97K
-OCID-Grasp [3] Rect. RGB-D 89 $\sim$ 7 75K 1763
+OCID-Grasp [53] Rect. RGB-D 89 $\sim$ 7 75K 1,763
 MetaGraspNet [54] 6D 3D 82 1-5K N/A 217K
-ACRONYM [12] 6D 3D 8872 2000 10.5M N/A
+ACRONYM [55] 6D 3D 8,872 2,000 10.5M N/A
 Grasp-Anything [56] Rect. RGB 3M 200 600M 1M
 ReasoingGrasp (ours) 6D RGB-D 64 $\sim$40K $\sim$99.3M 97K
-V-A Reasoning Instruction Generation
-As illustrated in Fig.
-3, we developed a unique and detailed method to create reasoning text instructions for our dataset by creating detailed object descriptions, automated generation with GPT-4, and manual review.
-The generated instructions enable robots to grasp based on nuanced, context-rich queries.
-V-A1 Description Creation
-The initial step in our instruction generation process involved creating detailed descriptions for each object and part, providing essential information such as functionalities, physical attributes, and typical uses.
-For instance, a pair of scissors was detailed as “A handheld cutting instrument with two crossing metal blades pivoted together, typically used for cutting paper or fabric.
-This particular pair has a black handle with yellow inner grips.” These descriptions served as the foundation for generating indirect questions and instructions.
-V-A2 Automated Generation with GPT-4
-The next step in instruction generation involved utilizing a customized variant of the GPT-4 model.
-This model was tasked with generating indirect questions and instructions for various objects and their parts.
-The output consists of 10 strings: 5 indirect questions that reference the object’s functions or design, and 5 indirect instructions that describe actions involving the object or its usage in certain tasks.
-These are crafted to be specific, relevant, and direct for grasping tasks.
-V-A3 Manual Review and Refinement
-Post-generation, each set of GPT-generated instructions underwent a manual review and refinement process.
-The manual intervention allowed us to fine-tune the language, enhancing the quality and applicability of the instructions for the targeted objects and their parts.
-This semi-automated approach balanced the efficiency of machine-generated content and judgment of human experts.
-V-B Part Annotation
-V-B1 Part Segmentation
-We manually segmented the part point clouds of each object using a specially developed annotation tool.
-The segmentation process was guided by an understanding of how humans typically interact with and use various parts of an object.
-For instance, a knife was divided into its blade and handle, reflecting the distinct functional roles of each part.
-V-B2 Grasping Pose Annotation
-Our dataset’s annotation process includes two key components: pose assignment and pose projection, for both 3D and 2D grasping poses, as shown in Fig. 4.
-Pose assignment involves aligning the grasping poses with the nearest part of the object.
-For each object’s segmented parts, we determine the grasping center of each point.
-These grasping poses are then assigned to the nearest part, ensuring an accurate association with the specific object part they correspond to.
-Pose projection transforms the grasping poses of parts into both 3D and 2D representations.
-This is achieved using camera matrices and object transformations.
-In the case of 3D, the original grasping poses are preserved in the spatial domain.
-For 2D projection, the 3D points of each segmented part are projected onto 2D images.
-This results in 2D masks that depict the spatial distribution of each part on the image plane.
-Grasping poses within these masks are then mapped to the corresponding parts in 2D.
-V-C Dataset Statistics
-In ensuring data quality, low-quality grasping poses were eliminated based on confidence evaluations and objects lacking semantic information were removed.
-In total, our enhanced dataset comprises an extensive collection of 1,730 instruction pairs, 64 objects, 109 segmented parts, and around 100 million associated grasping poses.
-Table I shows the statistics comparison of the previous grasping and visual reasoning datasets with ours.
-Unlike other grasping datasets, we offer detailed reasoning instructions and part-specific grasping information.
-VI Experiment
-In this section, we report the experiment results of the proposed method on the reasoning grasping dataset and real-world experiments.
-In the following experiments, we utilize both explicit instructions, such as “Pick up the <object>”, as well as implicit instructions that do not directly mention the grasping target.
-VI-A Implementation Details
-For the multi-modal LLM in the proposed model, we utilize the pre-trained LLaVA-7B-v0, which is derived from the large language model LLaMA-7B.
-As for the LoRA fine-tuning, we choose a rank $r=64$.
-During the training, we use the batch size of $8$, and the learning rate is set to $45e-$ 4, utilizing a cosine annealing schedule for adaptive learning rate adjustment.
-For the relative weight parameters in the loss function 1, we set $\lambda_{t}=1$ and $\lambda_{g}=1$ for the initial configuration, and starting from the third epoch, the parameter $\lambda_{t}$ is adjusted to 00.
-For the train-test-split, we use $90\%$ of the first 100 scenes in GraspNet-1 billion [13] for training and $10\%$ of the first 100 scenes for testing.
-We sampled 50k from the original LLaVA Instruct 150K dataset [24] to mix with 100k grasping data to maintain the visual reasoning capabilities.
-VI-B Baselines
-For the problem of reasoning grasping, we implement two baselines to compare with our proposed model.
-One baseline employs a CLIP text encoder [34] to extract textual features, while the other utilizes a modular approach incorporating the latest LLaVA model.
-VI-B1 CLIP $+$ GR-ConvNet Baseline
-In this baseline, we combine a CLIP text encoder with a CNN-based grasp detection model, GR-ConvNet [21].
-We first utilize the CLIP text encoder to extract the textual feature from input instructions, and this feature is then integrated into the hidden layers of GR-ConvNet.
-While CLIP can effectively extract language features relevant to visual content, it may fall short in capturing implicit reasoning compared to the VLM.
-VI-B2 LLaVa $\rightarrow$ GR-ConvNet Baseline
-This is a modular baseline that integrates the latest LLaVA-1.6-34b [25] with a pre-trained grasping detection model GR-ConvNet.
-As LLaVA cannot directly output the grasping pose, this baseline operates in two stages to obtain the grasping pose: first, LLaVA takes an image and an instruction as the input and is prompted to output a bounding box of the target object in the image.
-Then GR-ConvNet detects the optimal grasping pose within this specified bounding box.
-Note that LLaVA-1.6-34b has a much higher performance than LLaVa-v0-7b [24], which is the base model of our proposed reasoning grasping model.
-This baseline mirrors a similar process to LAN-grasp.
-VI-C Results
-VI-C1 Text Generation
-We first evaluate the precision of text generation by the fine-tuned multi-modal LLM in the reasoning module.
-Accurately generating special tokens and target names is crucial, since this identified grasping target will be utilized in the subsequent grasping module.
-The special tokens and target names act as intermediate states and enable the model to eventually generate a precise grasping pose from given instructions.
-We differentiate between explicit instructions, which directly mention the object’s name, and implicit instructions, which hint at the target object without stating its name.
-Note that the instructions used for testing were not seen during the training phase.
-Results, presented in Table II, show that our model can successfully interpret $92.52\%$ and $88.79\%$ of implicit instructions for object grasping and part grasping, respectively.
-This demonstrates our model’s good reasoning capabilities in understanding the grasping instructions.
-TABLE II: Text generation accuracy.
-Explicit Instruction Implicit Instruction
-Object Grasping 99.01% 92.52%
-Part Grasping 95.33% 88.79%
-TABLE III: Grasp prediction accuracy on the reasoning grasping dataset.
-Explicit Instruction Implicit Instruction Explicit Instruction Implicit Instruction
-Object Grasping Object Grasping Part Grasping Part Grasping
-Models R@1 R@2 R@3 R@1 R@2 R@3 R@1 R@2 R@3 R@1 R@2 R@3
-LLaVA $\rightarrow$ GR-ConvNet 26.80% 39.17% 44.66% 18.34% 23.76% 31.29% 15.38% 23.52% 37.93% 0.92% 3.84% 8.53%
-CLIP $+$ GR-ConvNet 50.40% 62.88% 66.38% 44.35% 57.15% 61.09% 43.78% 55.71% 59.61% 28.88% 42.34% 47.37%
-Reasoning Grasping 64.49% 84.11% 86.92% 63.55% 73.83% 77.57% 59.81% 70.01% 81.31% 61.68% 71.96% 83.18%
-LLaVA $\rightarrow$ GR-ConvNet (with depth) 31.88% 40.72% 45.95% 20.39% 25.77% 34.84% 20.11% 29.30% 36.63% 1.36% 2.77% 11.95%
-CLIP $+$ GR-ConvNet (with depth) 54.11% 66.50% 69.57% 45.89% 58.11% 62.50% 45.53% 59.51% 62.27% 33.06% 48.55% 52.25%
-Reasoning Grasping (with depth) 60.75% 71.95% 76.63% 57.94% 70.09% 77.46% 69.16% 72.90% 77.55% 64.48% 75.70% 80.37%
-VI-C2 Reasoning Grasping Result
-Table III shows the grasp prediction results on the reasoning grasping dataset.
-Performance is reported using the rectangle evaluation metric [18].
-A grasp pose is deemed valid if it fulfills the following two conditions: 1) The Intersection over Union (IoU) score between the predicted and ground truth rectangles exceeds $25\%$; 2) The angular deviation between the orientations of the predicted and ground truth rectangles is less than 30 degrees.
-We reported R@$k$, indicating the presence of valid grasps within the top-$k$ grasp predictions.
-Four distinct scenarios are evaluated, categorized by the combination of instruction types (explicit or implicit) and grasping targets (object or part).
-We assess both baseline models and our reasoning grasping model with and without depth input.
-Note that when depth information is utilized in the reasoning grasping model, it only serves as input for the grasping module and is not incorporated into the reasoning module, as shown in Fig.
-1.
-The results show that our reasoning grasping model outperforms the LLaVA $\rightarrow$ GR-ConvNet and CLIP $+$ GR-ConvNet baselines across all scenarios.
-The first baseline, LLaVA $\rightarrow$ GR-ConvNet, shows limited effectiveness, as the direct combining of two models fails to convey sufficient information for accurate grasp detection.
-Although the CLIP $+$ GR-ConvNet baseline exhibits some improvements, it still falls short, particularly in scenarios involving part grasping with implicit instructions.
-This indicates CLIP’s limited capability in processing implicit instructions.
-In contrast, our model excels in all scenarios, both with and without depth input, showcasing its superior ability to interpret implicit instructions and accurately generate the corresponding grasping poses.
-Note that some common failure cases are distinguishing objects with similar appearances, such as a shampoo bottle and a hair conditioner bottle, or identifying specific parts of complex objects like the uniformly colored toy camel.
-VI-C3 Visual Reasoning Ability
-Besides reasoning grasping, we also examine the visual reasoning capabilities of our model.
-The objective is to determine whether the model, after being fine-tuned, maintains the visual comprehension skills of the original LLaVA model.
-For a quantitative evaluation of performance, we employ GPT-4 to assess the quality of responses generated by our model.
-This is inspired by the evaluation methods in [24, 7].
-For each given image and question pair, both the original LLaVa model and our fine-tuned version provide answers related to the input image.
-We subsequently submit the questions, and answers from both LLaVA and our model, along with the ground-truth text descriptions, to a text-only GPT-4-based judge.
-This judge assesses the responses based on their helpfulness, relevance, accuracy, and level of detail in comparison to the ground-truth descriptions.
-GPT-4 then allocates an overall performance score ranging from 0 to 10, where higher scores denote superior performance.
-The image-question-answer sets used in this evaluation are randomly chosen from the LLaVa-Instruct-150k dataset [24].
-According to the judge of GPT-4, the original LLaVA-7B-v0 achieves $8.25/10$ when taking ground-truth textual descriptions as references.
-On the other hand, our model achieves $7.43/10$ under the same GPT-4 judge.
-This shows that our fine-tuned model still preserves a good portion of the reasoning ability of the original LLaVA.
-It also shows that the multi-round conversation ability is preserved, as the image-question-answer sets include sequences of multi-round conversations.
-VI-D Real World Experiments
-Figure 5: Real world experiments with implicit instructions.
-We conduct real-world experiments to evaluate the proposed reasoning grasping model and the CLIP $+$ GR-ConvNet baseline.
-A 7-DoF Franka Emika Panda equipped with a Franka Hand parallel gripper was utilized for grasping execution.
-To perceive the objects, one Azure Kinect was positioned in front of the robot.
-A cropped $480\times 480$ RGB image of the workspace was utilized as the input to the model.
-To execute the 2D grasps generated from the model, we always apply a top-down grasp with the height computed using the depth of the grasping point.
-In our experiments, we randomly place 4-8 objects on a tabletop and provide either explicit or implicit language instruction to the robot to grasp an object or a specific part, as shown in Fig.
-5.
-We select objects that are either identical or similar to those in the dataset.
-We deliberately excluded any objects or parts that are not infeasible for grasping, such as items too wide for the robot’s gripper.
-The performance is evaluated using three evaluation metrics: the correctness of generating special tokens and grasping target names, the accuracy of the output grasp pose, and the success in lifting the object.
-The results are reported in Table IV.
-Overall our proposed model can generate accurate special tokens and target names 39 out of 40 trials.
-This again shows the excellent reasoning capabilities of our model to interpret the various instructions.
-The failure case involved the model incorrectly identifying a screwdriver instead of scissors in response to a prompt requesting a tool to open a package box.
-We classified it as a failure solely because it did not match the specific pairing in our dataset, which serves as the basis for our evaluation metrics.
-This example underscores that even when the model’s response is technically incorrect, it is still reasonable.
-In terms of grasping performance, we assessed top-3 output grasp poses and manually selected the most feasible one for execution on the robot arm.
-We conducted 10 trials for each scenario, totaling 80 trials, to compare reasoning grasping with CLIP $+$ GR-ConvNet baseline.
-The CLIP $+$ GR-ConvNet baseline and our reasoning grasping model produced 15/40 and 25/40 accurate grasping poses, respectively, with the success rates for lifting objects are 9/40 and 16/40, respectively.
-The main reasons for accurate poses but fail to lift were due to the object slipping from the gripper or the object being too heavy.
-The experiment results demonstrate that our model outperforms the baseline in four distinct scenarios, showing its superior ability in reasoning grasping tasks.
-TABLE IV: Real-world experiment results.
-CLIP + GR-ConvNet Reasoning Grasping
-Object Grasping with Explicit Instruction
-Token Accuracy
-N/A 10/10
-Pose Accuracy
-5/10 7/10
-Execution Success
-2/10 5/10
-Object Grasping with Implicit Instruction
-Token Accuracy
-N/A 9/10
-Pose Accuracy
-3/10 5/10
-Execution Success
-2/10 3/10
-Part Grasping with Explicit Instruction
-Token Accuracy
-N/A 10/10
-Pose Accuracy
-5/10 7/10
-Execution Success
-3/10 4/10
-Part Grasping with Implicit Instruction
-Token Accuracy
-N/A 10/10
-Pose Accuracy
-2/10 6/10
-Execution Success
-2/10 4/10
-VII Conclusion
-In this paper, we introduce a novel task of reasoning grasping, where robots need to interpret and generate grasping poses based on implicit instructions.
-By leveraging the reasoning ability in a multi-modal large language model, our proposed model can understand indirect verbal commands and generate corresponding grasping poses.
-In addition, we also present the first dataset for reasoning grasping, derived from the GraspNet-1 billion dataset, featuring implicit human instructions alongside object and part grasping annotations.
-Experiment results demonstrate that our approach can effectively comprehend implicit instructions and accurately generate corresponding grasping poses.
-Overall, this work offers valuable insights bridging implicit human instructions and generating precise robot actions.
-While the proposed reasoning grasping model shows promising results, the model is limited when generating optimal grasping poses for novel objects and refining poses through conversational interactions.
-A possible reason is that the variety of objects with grasping annotations is still limited.
-For future work, the LLaVA-v0-7b model is not the state-of-the-art Visual Language Model.
-Other models such as LLaVA-1.6 [25] or GPT-4V [48] have shown much better performance and stronger ability in terms of visual understanding.
-Our model’s architecture, with its modular reasoning and grasping components, allows for flexibility in upgrading to more sophisticated models in the future.
-Appendix A Object and Part Statistics
-TABLE V: Selected Objects and Parts for the Training Split of the Dataset
+Appendix H Object and Part Statistics
+Table 5: Selected Objects and Parts for the Training Split of the Dataset
 Object Name Part Name
@@
 Scenes
-0 Cracker Box - 10971 4672 35
-1 Tomato Soup Can - 16625 8284 34
-5 Banana Stem, Flesh 1298 2941 25
-7 Red Mug Handle, Rim, Body 3363 2045 25
-8 Power Drill Handle, Body, Batterypack, Chuck 2445 4476 26
+0 Cracker Box - 10,971 4,672 35
+1 Tomato Soup Can - 16,625 8,284 34
+5 Banana Stem, Flesh 1,298 2,941 25
+7 Red Mug Handle, Rim, Body 3,363 2,045 25
+8 Power Drill
+Handle, Chuck,
+2,445 4,476 26
+Battery Pack, Body
 9 Scissor Handle, Blade 548 574 22
 11 Strawberry - 130 903 24
-14 Peach - 960 1922 25
-15 Pear - 948 2069 26
-17 Orange - 581 1466 32
-18 Knife Handle, Blade 1975 1526 29
-20 Red Screwdriver Handle, Shaft 1191 1910 29
+14 Peach - 960 1,922 25
+15 Pear - 948 2,069 26
+17 Orange - 581 1,466 32
+18 Knife Handle, Blade 1,975 1,526 29
+20 Red Screwdriver Handle, Shaft 1,191 1,910 29
 21 Racquetball - 443 573 6
-22 Blue Cup Rim, Body 3960 2770 29
-36 Daobao Wash Soap Cap, Body 2594 1744 25
-37 Nzskincare Mouth Rinse Cap, Body 1806 1786 24
-38 Daobao Sod Cap, Body 3256 2227 25
-40 Kispa Cleanser Cap, Body 931 1654 25
-41 Darlie Toothpaste Cap, Body 442 1181 25
+22 Blue Cup Rim, Body 3,960 2,770 29
+36 Daobao Wash Soap Cap, Body 2,594 1,744 25
+37 Nzskincare Mouth Rinse Cap, Body 1,806 1,786 24
+38 Daobao Sod Cap, Body 3,256 2,227 25
+40 Kispa Cleanser Cap, Body 931 1,654 25
+41 Darlie Toothpaste Cap, Body 442 1,181 25
 43 Baoke Marker Cap, Body 953 697 21
-44 Hosjam Toothpaste Pump Pump, Body 725 1958 22
+44 Hosjam Toothpaste Pump Pump, Body 725 1,958 22
 46 Dish Rim, Body 8 22 25
 48 Camel Legs, Head, Body 96 663 25
-51 Elephant Legs, Head, Body, Truck 218 1692 24
-52 Rhinocero Legs, Head, Body, Horn 152 1419 22
-57 Weiquan Chicken Bouillon Cap, Body 2297 2617 25
-58 Darlie Toothpaste Box - 5526 1646 25
+51 Elephant
+Legs, Head,
+218 1,692 24
+Body, Truck
+52 Rhinocero
+Legs, Head,
+152 1419 22
+Body, Horn
+57 Weiquan Chicken Bouillon Cap, Body 2,297 2,617 25
+58 Darlie Toothpaste Box - 5,526 1,646 25
 60 Black Mouse - 170 618 25
-61 Dabao Facewash Cap, Body 4683 2297 22
-62 Pantene Shampoo Cap, Body 502 1175 26
-63 Head Shoulders Supreme Cap, Body 777 1700 24
-66 Head Shoulders Care Cap, Body 1338 1269 24
-70 Tape - 4465 1859 25
-Total 33 54 76377 64355 100
-Appendix B 3D Object and Part Visualization
-Figure 6: 3D Visualization of Selected Objects and its Parts
-Appendix C Prompt for Instruction Generation
-Figure 7: Example Prompts used for Initial Reasoning Instruction Generation with GPT-4
+61 Dabao Facewash Cap, Body 4,683 2,297 22
+62 Pantene Shampoo Cap, Body 502 1,175 26
+63 Head Shoulders Supreme Cap, Body 777 1,700 24
+66 Head Shoulders Care Cap, Body 1,338 1,269 24
+70 Tape - 4,465 1,859 25
+Appendix I Hyperparameters
+For the multi-modal LLM in the proposed model, we utilize the pre-trained LLaVA-7B-v0, which is derived from the large language model LLaMA-7B.
+As for the LoRA fine-tuning, we choose a rank $r=64$.
+During the training, we use the batch size of $8$, and the learning rate is set to $45e-$ 4, utilizing a cosine annealing schedule for adaptive learning rate adjustment.
+For the relative weight parameters in the loss function 1, we set $\lambda_{t}=1$ and $\lambda_{g}=1$ for the initial configuration, and starting from the third epoch, the parameter $\lambda_{t}$ is adjusted to 00.
+For the train-test-split, we use $90\%$ of the first 100 scenes in GraspNet-1 billion  [19] for training and $10\%$ of the first 100 scenes for testing.
+We sampled 50k from the original LLaVA Instruct 150K dataset [12] to mix with 100k grasping data to maintain the visual reasoning capabilities.
+Appendix J Experiments on Visual Reasoning Ability.
+To evaluate the visual reasoning capabilities of our model, we employ a GPT-4 judge to assess the quality of responses generated by our model.
+For each given image and question pair, both the original LLaVa model and our fine-tuned version provide answers related to the input image.
+We subsequently submit the questions, and answers from both LLaVA and our model, along with the ground-truth text descriptions, to a text-only GPT-4-based judge.
+This judge assesses the responses based on their helpfulness, relevance, accuracy, and level of detail in comparison to the ground-truth descriptions.
+GPT-4 then allocates an overall performance score ranging from 0 to 10, where higher scores denote superior performance.
+The image-question-answer sets used in this evaluation are randomly chosen from the LLaVa-Instruct-150k dataset [12] with multi-round conversations.